### v1

# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

Abstract. The notion of probability plays a crucial role in quantum mechanics. It appears in quantum mechanics as the Born rule. In modern mathematics which describes quantum mechanics, however, probability theory means nothing other than measure theory, and therefore any operational characterization of the notion of probability is still missing in quantum mechanics. In our former works [K. Tadaki, arXiv:1804.10174], based on the toolkit of algorithmic randomness, we presented a refinement of the Born rule, called the principle of typicality, for specifying the property of results of measurements in an operational way. The Wigner’s friend paradox is a Gedankenexperiment regarding when and where the reduction of the state vector occurs in a chain of the measurements by several observers where the state of the consciousness of each observer is measured by the subsequent observer, except for the last observer in the chain. It is one of the central open questions in the measurement problem of quantum mechanics. In this paper, we make an analysis of the Wigner’s friend paradox in the framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox. In addition, we make a prediction, which is testable in principle, about its variant proposed by Deutsch.
Key words: the principle of typicality, algorithmic randomness, Born rule, probability interpretation, operational characterization, Martin-Löf randomness, many-worlds interpretation, reduction of the state vector, Schrödinger’s cat, Wigner’s friend, Deutsch’s thought experiment
1 Introduction
The notion of probability plays a crucial role in quantum mechanics. It appears in quantum mechanics as the so-called Born rule, i.e., the probability interpretation of the wave function [13, 53, 30]. In modern mathematics which describes quantum mechanics, however, probability theory means nothing other than measure theory, and therefore any operational characterization of the notion of probability is still missing in quantum mechanics. In this sense, the current form of quantum mechanics is considered to be imperfect as a physical theory which must stand on operational means.
In the work [37, 38, 41, 44, 48] we developed a theory of operational characterization of the notion of probability by algorithmic randomness for general discrete probability spaces, using the notion of Martin-Löf randomness with respect to Bernoulli measure [29] (and its extension over the Baire space). In the work we gave natural and equivalent operational characterizations of the basic notions of probability theory, such as the notions of conditional probability and the independence of events/random variables, in terms of the notion of Martin-Löf randomness with respect to Bernoulli measure. We then made applications of this framework to information theory and cryptography, as examples of the applications, in order to demonstrate the wide applicability of our framework to the general areas of science and technology. The crucial point is that in our theory the underlying discrete probability space is quite arbitrary, and thus we do not need to impose any computability restrictions on the underlying discrete probability space at all. See Tadaki [44] for the details of the theory for general finite probability spaces and its applications, and Tadaki [48] for ones for general discrete probability spaces whose sample spaces are countably infinite.
As a major application of our framework [37, 38, 41, 44] to basic science, in the work [36, 39, 40, 42, 43, 45, 46, 47] we presented a refinement of the Born rule, as an alternative rule to it, based on our theory of operational characterization of the notion of probability by algorithmic randomness, for the purpose of making quantum mechanics operationally perfect. Namely, we used the notion of Martin-Löf randomness with respect to Bernoulli measure to present the refinement of the Born rule, for specifying the property of the results of quantum measurements in an operational way. We then presented an operational refinement of the Born rule for mixed states, as an alternative rule to it, based on Martin-Löf randomness with respect to Bernoulli measure. In particular, we gave a precise definition for the notion of mixed state. We then showed that all of the refined rules of the Born rule for both pure states and mixed states can be derived from a single postulate, called the principle of typicality, in a unified manner. We did this from the point of view of the many-worlds interpretation of quantum mechanics [22]. We then made an application of our framework to the BB84 quantum key distribution protocol in order to demonstrate how properly our framework works in practical problems in quantum mechanics, based on the principle of typicality. See the work [47] for the details of our framework of quantum mechanics based on the principle of typicality.
In the subsequent work [49], we made an application of our framework based on the principle of typicality to the argument of local realism versus quantum mechanics for refining it, in order to further demonstrate how properly our framework works in practical problems in quantum mechanics.
First, in the work [49] we refined and reformulated the argument of Bell’s inequality versus quantum mechanics [1], by means of algorithmic randomness. On the one hand, we refined and reformulated the assumptions of local realism to lead to Bell’s inequality, in terms of the theory of operational characterization of the notion of probability by algorithmic randomness, developed by the work [37, 38, 44] as mentioned above. On the other hand, we refined and reformulated the corresponding argument of quantum mechanics to violate Bell’s inequality, based on the principle of typicality.
Second, in the work [49] we refined and reformulated an entirely different type of the variant of the argument of Bell’s inequality versus quantum mechanics, i.e., the argument over the GHZ experiment [19, 20, 28], in the same manner as above. In the case of the GHZ experiment, we made the ‘two-sided’ refinements regarding both the predictability (by quantum mechanics) and unexplainability (by local realism) of the perfect correlations of measurement results over three parties for a system of three spin-$1/2$ particles, in terms of algorithmic randomness.
On the other hand, in the work [50], we made an application of our framework based on the principle of typicality to the theory of quantum error-correction for refining it, in order to further demonstrate how properly our framework works in practical problems in quantum mechanics.
The general theory of quantum error-correcting codes is introduced and developed by Ekert and Macchiavello [15], Bennett, et al. [2], and Knill and Laflamme [25]. In the general theory, the quantum error-correction conditions give necessary and sufficient conditions for quantum error-correction to be possible. To show the sufficiency of the quantum error-correction conditions involves the construction of an error-correction operation. In the work [50], we refined and reformulated the whole process consisting of the noise process and the error-correction operation in terms of our framework based on the principle of typicality, in order to operationally show that quantum error-correction is possible if the quantum error-correction conditions hold.
The Calderbank-Shor-Steane codes (CSS codes, for short) is a large class of quantum error-correcting codes, introduced and developed by Calderbank and Shor [6] and Steane [34]. In the work [50], we refined and reformulated the whole process consisting of the noise process and the error-correction operation for CSS codes, based on the principle of typicality, in order to show in an operational manner that quantum error-correction by CSS codes is possible.
Around 2000, there were comprehensive attempts to extend the algorithmic randomness notions to the quantum regime. Namely, Berthiaume, van Dam, and Laplante [3], Vitányi [51], and Gács [16] independently introduced the so-called quantum Kolmogorov complexity from each standpoint, in order to define the algorithmic information content of a given individual quantum state of a finite dimensional quantum system. Later on, Tadaki [35] introduced in 2004 the notion of $\hat{\Omega}$ by means of modifying the work of Gács [16]. This $\hat{\Omega}$ is an extension of Chaitin’s halting probability $\Omega$ [9] to a measurement operator in an infinite dimensional quantum system. Actually, $\hat{\Omega}$ is a bounded Hermitian operator on a Hilbert space of infinite dimension, and for every computable state $|\Psi\rangle$ it holds that the inner-product $\langle\Psi|\hat{\Omega}|\Psi\rangle$, which has a meaning of “probability” in quantum mechanics, is a Martin-Löf random recursively enumerable real and therefore can be represented as a Chaitin’s $\Omega$. These previous works are about the extension of algorithmic randomness notions to the quantum realm.
In contrast, in this paper as well as our former works [36, 39, 40, 42, 43, 45, 46, 47, 49, 50], we do not extend the algorithmic randomness notions themselves, while keeping them in their original forms, but we identify in the mathematical framework of quantum mechanics the basic idea of Martin-Löf randomness that the random infinite sequences are precisely sequences which are not contained in any effective null set. See Martin-Löf [29], Nies [31], Downey and Hirschfeldt [14], and Brattka, Miller, and Nies [5] for this basic idea of Martin-Löf randomness. In our works, we do this identification, in particular, from the aspect of the many-worlds interpretation of quantum mechanics [22], leading to the principle of typicality.
1.1 Contributions of the paper
The Wigner’s friend paradox [54] is a Gedankenexperiment regarding when and where the reduction of the state vector occurs in a chain of the measurements by several observers where the state of the consciousness of each observer is measured by the subsequent observer, except for the last observer in the chain. The Wigner’s friend paradox is one of the central open questions in the measurement problem of quantum mechanics.
In this paper, we extend the framework of the principle of typicality, which was introduced by Tadaki [42, 43, 45, 46, 47], so that it can be applicable to the situation, such as the Wigner’s friend paradox, where apparatuses perform measurements over other apparatuses as well as over a normal quantum system only being measured. For that purpose, we introduce postulates, Postulates 6 and 7 below, for determining the states of apparatuses in early stages among all apparatuses constitute a measurement process. We make an analysis of the Wigner’s friend paradox within this extended framework of quantum mechanics based on the principle of typicality. We then draw common sense conclusions about the Wigner’s friend paradox.
Deutsch’s thought experiment [11] is a variant of the Wigner’s friend paradox, which can, in principle, verify the effect of the consciousness of observer on the reduction of the state vector. In this paper, we make an analysis of the Deutsch’s thought experiment comprehensively within the extended framework of the principle of typicality. Based on this, we make a prediction which is testable in principle.
Based on the extended framework of the principle of typicality, we can analyze still more complicated situations than the Wigner’s friend paradox or Deutsch’s thought experiment. As such a example, we introduce the combination of the Wigner’s friend paradox and Deutsch’s thought experiment, called the Wigner-Deutsch collaboration. We make an analysis of it comprehensively in this paper.
1.2 Organization of the paper
The paper is organized as follows. We begin in Section 2 with some mathematical preliminaries, in particular, about measure theory and Martin-Löf randomness with respect to an arbitrary probability measure. We then review the notion of Martin-Löf randomness with respect to an arbitrary Bernoulli measure, called the Martin-Löf $P$-randomness in this paper, in Section 3. In Section 4 we summarize the theorems and notions on Martin-Löf $P$-randomness from Tadaki [37, 38, 44], which are need to establish the contributions of this paper presented in the later sections. In Section 5 we review the central postulates of the conventional quantum mechanics according to Nielsen and Chuang [30]. In Section 6 we review the framework of the principle of typicality, which was introduced by Tadaki [42, 43, 45, 46, 47].
In Section 7 we extend the framework of the principle of typicality so that it is applicable to the situation where apparatuses perform measurements over other apparatuses, and we introduce postulates for determining the states of apparatuses in early stages of a measurement process. In Section 8 we review the Wigner’s friend paradox [54]. Based on this, in Section 9, we make an analysis of the Wigner’s friend paradox within the extended framework of the principle of typicality, developed in Section 7. On the other hand, in Section 10, we review Deutsch’s thought experiment [11]. Based on this, in Section 11, we make an analysis of Deutsch’s thought experiment within the extended framework of the principle of typicality. In Section 12 we make an analysis of Deutsch’s thought experiment where the ‘Friend’ is treated as a mere quantum system and not as a measurement apparatus, within the original framework of the principle of typicality, reviewed in Section 6.
In Section 13 we introduce the Wigner-Deutsch collaboration, which is a combination of the Wigner’s friend paradox and Deutsch’s thought experiment, and we make an analysis of it in the terminology of the conventional quantum mechanics. In Section 14 we make an analysis of the Wigner-Deutsch collaboration within the extended framework of the principle of typicality. Based on the extended framework developed in Section 7, we can analyze such a complicated situation. In Section 15 we make an analysis of the Wigner-Deutsch collaboration where the ‘Friend’ is treated as a mere quantum system and not as a measurement apparatus, within the extended framework of the principle of typicality.
2 Mathematical preliminaries
2.1 Basic notation and definitions
We start with some notation about numbers and strings which will be used in this paper. We denote the cardinality of $S$ by #S\#S for any finite set $S$. $\mathbb{N}=\left\{0,1,2,3,\dotsc\right\}$ is the set of natural numbers, and $\mathbb{N}^{+}$ is the set of positive integers.
An alphabet is a non-empty finite set. Let $\Omega$ be an arbitrary alphabet throughout the rest of this subsection. A finite string over $\Omega$ is a finite sequence of elements from the alphabet $\Omega$. We use $\Omega^{*}$ to denote the set of all finite strings over $\Omega$, which contains the empty string denoted by $\lambda$. For any $\sigma\in\Omega^{*}$, $\left\lvert\sigma\right\rvert$ is the length of $\sigma$. Therefore $\left\lvert\lambda\right\rvert=0$. A subset $S$ of $\Omega^{*}$ is called prefix-free if no string in $S$ is a prefix of another string in $S$.
An infinite sequence over $\Omega$ is an infinite sequence of elements from the alphabet $\Omega$, where the sequence is infinite to the right but finite to the left. We use $\Omega^{\infty}$ to denote the set of all infinite sequences over $\Omega$.
Let $\alpha\in\Omega^{\infty}$. For any $n\in\mathbb{N}$ we denote by $\alpha\!\!\restriction_{n}\in\Omega^{*}$ the first $n$ elements in the infinite sequence $\alpha$, and for any $n\in\mathbb{N}^{+}$ we denote by $\alpha(n)$ the $n$th element in $\alpha$. Thus, for example, $\alpha\!\!\restriction_{4}=\alpha(1)\alpha(2)\alpha(3)\alpha(4)$, and $\alpha\!\!\restriction_{0}=\lambda$.
For any $S\subset\Omega^{*}$, the set $\{\alpha\in\Omega^{\infty}\mid\exists\,n\in\mathbb{N}\;\alpha\!\!\restriction_{n}\in S\}$ is denoted by $\left[S\right]^{\prec}$. Note that (i) $\left[S\right]^{\prec}\subset\left[T\right]^{\prec}$ for every $S\subset T\subset\Omega^{*}$, and (ii) for every set $S\subset\Omega^{*}$ there exists a prefix-free set $P\subset\Omega^{*}$ such that $\left[S\right]^{\prec}=\left[P\right]^{\prec}$. For any $\sigma\in\Omega^{*}$, we denote by $\left[\sigma\right]^{\prec}$ the set $\left[\{\sigma\}\right]^{\prec}$, i.e., the set of all infinite sequences over $\Omega$ extending $\sigma$. Therefore $\left[\lambda\right]^{\prec}=\Omega^{\infty}$.
Let $\mathcal{H}$ be a complex Hilbert space. A projector $P$ in $\mathcal{H}$ is a non-zero Hermitian bounded operator in $\mathcal{H}$ such that $P^{2}=P$. A collection $\{P_{a}\}_{a\in\Theta}$ of projectors in $\mathcal{H}$ is called a projection-valued measure (PVM, for short) if $\Theta$ is an alphabet and the following hold:
  1. (i)
    $P_{a}P_{b}=\delta_{a,b}P_{a}$ for every $a,b\in\Theta$.
  2. (ii)
    $\sum_{a\in\Theta}P_{a}=I$, where $I$ denotes the identity operator in $\mathcal{H}$.
2.2 Martin-Löf randomness with respect to an arbitrary probability measure
We introduce the notion of Martin-Löf randomness with respect to an arbitrary probability measure. For that purpose, we briefly review measure theory according to Nies [31, Section 1.9]. See also Billingsley [4] for measure theory in general. We here summurize the description of Tadaki [44, Section 2.2], which is based on Nies [31, Section 1.9].
Let $\Omega$ be an arbitrary alphabet. A probability measure representation over $\Omega$ is a function $r\colon\Omega^{*}\to[0,1]$ such that (i) $r(\lambda)=1$ and (ii) for every $\sigma\in\Omega^{*}$ it holds that
$r(\sigma)=\sum_{a\in\Omega}r(\sigma a).$
A probability measure representation over $\Omega$ induces a probability measure on $\Omega^{\infty}$ in the following manner.
A class $\mathcal{F}$ of subsets of $\Omega^{\infty}$ is called a $\sigma$-field on $\Omega^{\infty}$ if $\mathcal{F}$ includes $\Omega^{\infty}$, is closed under complements, and is closed under the formation of countable unions. A subset $\mathcal{R}$ of $\Omega^{\infty}$ is called open if $\mathcal{R}=\left[S\right]^{\prec}$ for some $S\subset\Omega^{*}$. The Borel class $\mathcal{B}_{\Omega}$ is the $\sigma$-field generated by all open sets on $\Omega^{\infty}$. Namely, the Borel class $\mathcal{B}_{\Omega}$ is defined as the intersection of all the $\sigma$-fields on $\Omega^{\infty}$ containing all open sets on $\Omega^{\infty}$. A real-valued function $\mu$ defined on the Borel class $\mathcal{B}_{\Omega}$ is called a probability measure on $\Omega^{\infty}$ if the following conditions hold:
  1. (i)
    $\mu\left(\emptyset\right)=0$ and $\mu\left(\Omega^{\infty}\right)=1$;
  2. (ii)
    $\mu\left(\bigcup_{i}\mathcal{D}_{i}\right)=\sum_{i}\mu\left(\mathcal{D}_{i}\right)$ for every sequence $\{\mathcal{D}_{i}\}_{i\in\mathbb{N}}$ of sets in $\mathcal{B}_{\Omega}$ such that $\mathcal{D}_{i}\cap\mathcal{D}_{i}=\emptyset$ for all $i\neq j$.
Then we can show that for every probability measure representation $r$ over $\Omega$ there exists a unique probability measure $\mu$ on $\Omega^{\infty}$ such that
$\mu\left(\left[\sigma\right]^{\prec}\right)=r(\sigma)$ (1)
for every $\sigma\in\Omega^{*}$. This unique probability measure $\mu$ is called the probability measure induced by the probability measure representation $r$, denoted $\mu_{r}$.
Now, we introduce the notion of Martin-Löf randomness [29] in a general setting, as follows.
Definition 2.1 (Martin-Löf randomness with respect to a probability measure).
Let $\Omega$ be an alphabet, and let $\mu$ be a probability measure on $\Omega^{\infty}$. A subset $\mathcal{C}$ of $\mathbb{N}^{+}\times\Omega^{*}$ is called a Martin-Löf test with respect to $\mu$ if $\mathcal{C}$ is a recursively enumerable set and for every $n\in\mathbb{N}^{+}$ it holds that $\mathcal{C}_{n}$ is a prefix-free subset of $\Omega^{*}$ and
$\mu\left(\left[\mathcal{C}_{n}\right]^{\prec}\right)<2^{-n},$ (2)
where $\mathcal{C}_{n}$ denotes the set $\left\{\,\sigma\mid(n,\sigma)\in\mathcal{C}\,\right\}$.
For any $\alpha\in\Omega^{\infty}$, we say that $\alpha$ is Martin-Löf random with respect to $\mu$ if
$\alpha\notin\bigcap_{n=1}^{\infty}\left[\mathcal{C}_{n}\right]^{\prec}$
for every Martin-Löf test $\mathcal{C}$ with respect to $\mu$.∎
[…]
The principle of typicality, Postulate 5 below, is stated by means of the notion of Martin-Löf randomness with respect to an arbitrary probability measure introduced in the preceding section. However, in many situations of the applications of the principle of typicality, such as in a contribution of this paper described in Sections 7, 9, 11, 12, 14, and 15, a more restricted notion is used where the probability measure is chosen to be a Bernoulli measure. Namely, the notion of Martin-Löf randomness with respect to a Bernoulli measure is used in many situations of the applications of the principle of typicality. Thus, in order to introduce this notion, we first review the notions of finite probability space and Bernoulli measure.
Definition 3.1 (Finite probability space).
Let $\Omega$ be an alphabet. A finite probability space on $\Omega$ is a function $P\colon\Omega\to[0,1]$ such that (i) $P(a)\geq 0$ for every $a\in\Omega$, and (ii) $\sum_{a\in\Omega}P(a)=1$. The set of all finite probability spaces on $\Omega$ is denoted by $\mathbb{P}(\Omega)$.
Let $P\in\mathbb{P}(\Omega)$. The set $\Omega$ is called the sample space of $P$, and elements of $\Omega$ are called sample points or elementary events of $P$. For each $A\subset\Omega$, we define $P(A)$ by
$P(A):=\sum_{a\in A}P(a).$
A subset of $\Omega$ is called an event on $P$, and $P(A)$ is called the probability of $A$ for every event $A$ on $P$. ∎
Let $\Omega$ be an alphabet, and let $P\in\mathbb{P}(\Omega)$. For each $\sigma\in\Omega^{*}$, we use $P(\sigma)$ to denote
$P(\sigma_{1})P(\sigma_{2})\dots P(\sigma_{n})$
where $\sigma=\sigma_{1}\sigma_{2}\dots\sigma_{n}$ with $\sigma_{i}\in\Omega$. Therefore $P(\lambda)=1$, in particular. For each subset $S$ of $\Omega^{*}$, we use $P(S)$ to denote
$\sum_{\sigma\in S}P(\sigma).$
Therefore $P(\emptyset)=0$, in particular.
Consider a function $r\colon\Omega^{*}\to[0,1]$ such that $r(\sigma)=P(\sigma)$ for every $\sigma\in\Omega^{*}$. It is then easy to see that the function $r$ is a probability measure representation over $\Omega$. The probability measure $\mu_{r}$ induced by the probability measure representation $r$ is called the Bernoulli measure on $\Omega^{\infty}$ (induced by the finite probability space $P$), denoted $\lambda_{P}$. The Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ satisfies that
$\lambda_{P}\left(\left[\sigma\right]^{\prec}\right)=P(\sigma)$ (3)
for every $\sigma\in\Omega^{*}$, due to (1).
The notion of Martin-Löf randomness with respect to a Bernoulli measure is defined as follows. We call it the Martin-Löf $P$-randomness, since it depends on a finite probability space $P$. This notion was, in essence, introduced by Martin-Löf [29], as well as the notion of Martin-Löf randomness with respect to Lebesgue measure.
Definition 3.2 (Martin-Löf $P$-randomness, Martin-Löf [29]).
Let $\Omega$ be an alphabet, and let $P\in\mathbb{P}(\Omega)$. For any $\alpha\in\Omega^{\infty}$, we say that $\alpha$ is Martin-Löf $P$-random if $\alpha$ is Martin-Löf random with respect to $\lambda_{P}$.∎
Martin-Löf random infinite binary sequences [29] are pircisely Martin-Löf $U$-random infinite binary sequences where $U$ is a finite probability space on $\{0,1\}$ such that $U(0)=U(1)=1/2$.
[…]
In order to obtain the results in this paper, we need several theorems on Martin-Löf $P$-randomness from Tadaki [37, 38, 44]. Originally, these theorems played a key part in developing the theory of operational characterization of the notion of probability based on Martin-Löf $P$-randomness in Tadaki [37, 38, 44]. We enumerate these theorems in this section.
4.1 Marginal distribution
First, the following theorem states that Martin-Löf $P$-randomness is closed under the formation of marginal distribution. See Tadaki [47, Theorem 9.4] for the proof.
Theorem 4.1 (Tadaki [47]).
Let $\Omega$ and $\Theta$ be alphabets. Let $P\in\mathbb{P}(\Omega\times\Theta)$, and let $\alpha\in(\Omega\times\Theta)^{\infty}$. Suppose that $\beta$ is an infinite sequence over $\Omega$ obtained from $\alpha$ by replacing each element $(m,l)$ in $\alpha$ by $m$. If $\alpha$ is Martin-Löf $P$-random then $\beta$ is Martin-Löf $Q$-random, where $Q\in\mathbb{P}(\Omega)$ such that
$Q(m):=\sum_{l\in\Theta}P(m,l)$
for every $m\in\Omega$. ∎
Note that in Theorems 4.1 the underlying finite probability space $P$ on $\Omega\times\Theta$ is quite arbitrary, and thus we do not impose any computability restrictions on $P$ at all, in particular.
4.2 The law of large numbers
The following theorem shows that the law of large numbers holds for an arbitrary Martin-Löf $P$-randomness infinite sequence. For the proof of Theorem 4.2, see Tadaki [44, Theorem 14].
Theorem 4.2 (The law of large numbers, Tadaki [37]).
Let $\Omega$ be an alphabet, and let $P\in\mathbb{P}(\Omega)$. For every $\alpha\in\Omega^{\infty}$, if $\alpha$ is Martin-Löf $P$-random then for every $a\in\Omega$ it holds that
$\lim_{n\to\infty}\frac{N_{a}(\alpha\!\!\restriction_{n})}{n}=P(a),$
where $N_{a}(\sigma)$ denotes the number of the occurrences of $a$ in $\sigma$ for every $a\in\Omega$ and every $\sigma\in\Omega^{*}$. ∎
Note that in Theorem 4.2 the underlying finite probability space $P$ is quite arbitrary, and thus we do not impose any computability restrictions on $P$ at all, in particular.
4.3 Non-occurrence of an elementary event with probability zero
From various aspects, Tadaki [44, Section 5.1] demonstrated the fact that an elementary event with probability zero never occurs in the conventional quantum mechanics. This fact corresponds to Theorem 4.3 below in the context of the theory of operational characterization of the notion of probability, which was developed by Tadaki [37, 38, 44] based on Martin-Löf $P$-randomness.
Theorem 4.3.
Let $\Omega$ be an alphabet, and let $P\in\mathbb{P}(\Omega)$. Let $a\in\Omega$. Suppose that $\alpha$ is Martin-Löf $P$-random and $P(a)=0$. Then $\alpha$ does not contain $a$. ∎
This result for Martin-Löf $P$-random was, in essence, pointed out by Martin-Löf [29]. Note that we do not impose any computability restrictions on the underlying finite probability space $P$ at all in Theorem 4.3. For the proof of Theorem 4.3, see Tadaki [44, Theorem 13]. The following corollary is immediate from Theorem 4.3.
Corollary 4.4.
Let $\Omega$ be an alphabet, and let $P\in\mathbb{P}(\Omega)$. For every $\alpha\in\Omega^{\infty}$, if $\alpha$ is Martin-Löf $P$-random then $P(\alpha(n))>0$ for every $n\in\mathbb{N}^{+}$. ∎
5 Postulates of quantum mechanics
In this section, we review the central postulates of (the conventional) quantum mechanics. For simplicity, in this paper we consider the postulates of quantum mechanics for a finite-dimensional quantum system, i.e., a quantum system whose state space is a finite-dimensional Hilbert space. Nielsen and Chuang [30] treat thoroughly the postulates of (the conventional) quantum mechanics in the finite-dimensional case, as a textbook of the field of quantum computation and quantum information in which such a case is typical. In this paper we refer to the postulates of the conventional quantum mechanics in the form presented in Nielsen and Chuang [30, Section 2.2].
The first postulate of quantum mechanics is about state space and state vector.
Postulate 1 (State space and state vector).
Associated to any isolated physical system is a complex vector space with inner product (that is, a Hilbert space) known as the state space of the system. The system is completely described by its state vector, which is a unit vector in the system’s state space. ∎
The second postulate of quantum mechanics is about the composition of systems.
Postulate 2 (Composition of systems).
The state space of a composite physical system is the tensor product of the state spaces of the component physical systems. Moreover, if we have systems numbered $1$ through $n$, and system number $i$ is prepared in the state $|\Psi_{i}\rangle$, then the joint state of the total system is $|\Psi_{1}\rangle\otimes|\Psi_{2}\rangle\otimes\dots\otimes|\Psi_{n}\rangle$. ∎
The third postulate of quantum mechanics is about the time-evolution of closed quantum systems.
Postulate 3 (Unitary time-evolution).
The evolution of a closed quantum system is described by a unitary transformation. That is, the state $|\Psi_{1}\rangle$ of the system at time $t_{1}$ is related to the state $|\Psi_{2}\rangle$ of the system at time $t_{2}$ by a unitary operator $U$, which depends only on the times $t_{1}$ and $t_{2}$, in such a way that $|\Psi_{2}\rangle=U|\Psi_{1}\rangle$. ∎
The forth postulate of quantum mechanics is about measurements on quantum systems.
Postulate 4 (Quantum measurement).
A quantum measurement is described by a collection $\{M_{m}\}_{m\in\Omega}$ of measurement operators which satisfy the completeness equation,
$\sum_{m\in\Omega}M_{m}^{\dagger}M_{m}=I.$
These are operators acting on the state space of the system being measured.
  1. (i)
    The set of possible outcomes of the measurement equals the finite set $\Omega$. If the state of the quantum system is $|\Psi\rangle$ immediately before the measurement then the probability that result $m$ occurs is given by
    $\langle\Psi|M_{m}^{\dagger}M_{m}|\Psi\rangle.$
  2. (ii)
    Given that outcome $m$ occurred, the state of the quantum system immediately after the measurement is
    $\frac{M_{m}|\Psi\rangle}{\sqrt{\langle\Psi|M_{m}^{\dagger}M_{m}|\Psi\rangle}}.$
Postulate 4 (i) is the so-called Born rule, i.e, the probability interpretation of the wave function, while Postulate 4 (ii) is called the projection hypothesis. Thus, the Born rule, Postulate 4 (i), uses the notion of probability. However, the operational characterization of the notion of probability is not given in the Born rule, and therefore the relation of its statement to a specific infinite sequence of outcomes of quantum measurements which are being generated by an infinitely repeated measurements is unclear. Tadaki [36, 39, 40, 42, 43, 45, 46, 47] fixed this point.
In this paper as well as our former works [36, 39, 40, 42, 43, 45, 46, 47, 49, 50], we keep Postulates 1, 2, and 3 in their original forms without any modifications. The principle of typicality, Postulate 5 below, is proposed as a refinement of Postulate 4 for replacing it, based on the notion of Martin-Löf randomness with respect to a probability measure.
6 The principle of typicality: Our world is typical
Everett [22] introduced the many-worlds interpretation of quantum mechanics (MWI, for short) in 1957. Actually, his MWI was more than just an interpretation of quantum mechanics. It aimed to derive Postulate 4 (i), the Born rule, from the remaining postulates, i.e., from Postulates 1, 2, and 3 above. In this sense, Everett [22] proposed his MWI as a “metatheory” of quantum mechanics. The point is that in MWI the measurement process is fully treated as the interaction between a system being measured and an apparatus measuring it, based only on Postulates 1, 2, and 3. Then his MWI tried to derive Postulate 4 (i) in such a setting. However, his attempt does not seem successful. One of the reasons is thought to be its mathematical vagueness.
Tadaki [42] reformulated the original framework of MWI by Everett [22] in a form of mathematical rigor from a modern point of view. The point of this rigorous treatment of the framework of MWI is the use of the notion of probability measure representation and its induction of probability measure, as presented in Section 2.2. Tadaki [42] then introduced the principle of typicality in this rigorous framework. The principle of typicality is an operational refinement of Postulate 4 based on the notion of Martin-Löf randomness with respect to a probability measure given in Definition 2.1. In what follows, we review this rigorous framework of MWI based the principle of typicality. See Tadaki [47] for the details of this framework of MWI and its validity.
Now, according to Tadaki [47], let us introduce the setting of MWI in terms of our terminology in a form of mathematical rigor. Let $\mathcal{S}$ be an arbitrary quantum system with state space $\mathcal{H}$ of finite dimension. Consider a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$ satisfying the completeness equation,
$\sum_{m\in\Omega}M_{m}^{\dagger}M_{m}=I.$ (4)
Here, $\Omega$ is the set of all possible outcomes of the measurement. [^1] The set $\Omega$ is finite, and therefore it is an alphabet. Let $\mathcal{A}$ be an apparatus performing the measurement described by $\{M_{m}\}_{m\in\Omega}$, which is a quantum system with state space $\overline{\mathcal{H}}$. According to Postulates 1, 2, and 3, the measurement process of the measurement described by the measurement operators $\{M_{m}\}_{m\in\Omega}$ is described by a unitary operator $U$ such that
$U(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)=\sum_{m\in\Omega}(M_{m}|\Psi\rangle)\otimes|\Phi[m]\rangle$ (5)
for every $|\Psi\rangle\in\mathcal{H}$. This is due to von Neumann [53, Section VI.3] in the case where the measurement operators $\{M_{m}\}_{m\in\Omega}$ form a PVM such that $\operatorname{rank}M_{m}=1$ for every $m\in\Omega$. For the derivation of (5) in the general case, see Tadaki [47, Section 7.1]. Actually, $U$ describes the interaction between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$. The vector $|\Phi_{\mathrm{init}}\rangle\in\overline{\mathcal{H}}$ is the initial state of the apparatus $\mathcal{A}$, and $|\Phi[m]\rangle\in\overline{\mathcal{H}}$ is a final state of the apparatus $\mathcal{A}$ for each $m\in\Omega$, with $\langle\Phi[m]|\Phi[m^{\prime}]\rangle=\delta_{m,m^{\prime}}$. For every $m\in\Omega$, the state $|\Phi[m]\rangle$ indicates that the apparatus $\mathcal{A}$ records the value $m$ as the measurement outcome. By the unitary interaction (5) as a measurement process, a correlation (i.e., entanglement) is generated between the system $\mathcal{S}$ and the apparatus $\mathcal{A}$.
In the framework of MWI, we consider countably infinite copies of the system $\mathcal{S}$, and consider a countably infinite repetition of the measurements described by the identical measurement operators $\{M_{m}\}_{m\in\Omega}$ performed over each of such copies in a sequential order, where each of the measurements is described by the unitary time-evolution (5). As repetitions of the measurement progressed, correlations between the systems and the apparatuses are being generated in sequence in the superposition of the total system consisting of the systems and the apparatuses. The detail is described as follows.
For each $n\in\mathbb{N}^{+}$, let $\mathcal{S}_{n}$ be the $n$th copy of the system $\mathcal{S}$ and $\mathcal{A}_{n}$ the $n$th copy of the apparatus $\mathcal{A}$. Each $\mathcal{S}_{n}$ is prepared in a state $|\Psi_{n}\rangle$ while all $\mathcal{A}_{n}$ are prepared in an identical state $|\Phi_{\mathrm{init}}\rangle$. The measurement described by the measurement operators $\{M_{m}\}_{m\in\Omega}$ is performed over each $\mathcal{S}_{n}$ one by one in the increasing order of $n$, by interacting each $\mathcal{S}_{n}$ with $\mathcal{A}_{n}$ according to the unitary time-evolution (5). For each $n\in\mathbb{N}^{+}$, let $\mathcal{H}_{n}$ be the state space of the total system consisting of the first $n$ copies $\mathcal{S}_{1},\mathcal{A}_{1},\mathcal{S}_{2},\mathcal{A}_{2},\dots,\mathcal{S}_{n},\mathcal{A}_{n}$ of the system $\mathcal{S}$ and the apparatus $\mathcal{A}$. These successive interactions between the copies of the system $\mathcal{S}$ and the apparatus $\mathcal{A}$ as measurement processes proceed in the following manner:
The starting state of the total system, which consists of $\mathcal{S}_{1}$ and $\mathcal{A}_{1}$, is $|\Psi_{1}\rangle\otimes|\Phi_{\mathrm{init}}\rangle\in\mathcal{H}_{1}$. Immediately after the measurement described by $\{M_{m}\}_{m\in\Omega}$ over $\mathcal{S}_{1}$, the total system results in the state
$\sum_{m_{1}\in\Omega}(M_{m_{1}}|\Psi_{1}\rangle)\otimes|\Phi[m_{1}]\rangle\in\mathcal{H}_{1}$
by the interaction (5) as a measurement process. In general, immediately before performing the measurement described by $\{M_{m}\}_{m\in\Omega}$ over $\mathcal{S}_{n}$, the state of the total system, which consists of $\mathcal{S}_{1},\mathcal{A}_{1},\mathcal{S}_{2},\mathcal{A}_{2},\dots,\mathcal{S}_{n},\mathcal{A}_{n}$, is
$\sum_{m_{1},\dots,m_{n-1}\in\Omega}(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n-1}}|\Psi_{n-1}\rangle)\otimes|\Psi_{n}\rangle\otimes|\Phi[m_{1}]\rangle\otimes\dots\otimes|\Phi[m_{n-1}]\rangle\otimes|\Phi_{\mathrm{init}}\rangle$
in $\mathcal{H}_{n}$, where $|\Psi_{n}\rangle$ is the initial state of $\mathcal{S}_{n}$ and $|\Phi_{\mathrm{init}}\rangle$ is the initial state of $\mathcal{A}_{n}$. Immediately after the measurement described by $\{M_{m}\}_{m\in\Omega}$ over $\mathcal{S}_{n}$, the total system results in the state
$\sum_{m_{1},\dots,m_{n}\in\Omega}(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}]\rangle\otimes\dots\otimes|\Phi[m_{n}]\rangle$
$=$ $\sum_{m_{1},\dots,m_{n}\in\Omega}(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}\dots m_{n}]\rangle$ (6)
in $\mathcal{H}_{n}$, by the interaction (5) as a measurement process between the system $\mathcal{S}_{n}$ prepared in the state $|\Psi_{n}\rangle$ and the apparatus $\mathcal{A}_{n}$ prepared in the state $|\Phi_{\mathrm{init}}\rangle$. The vector $|\Phi[m_{1}\dots m_{n}]\rangle$ denotes the vector $|\Phi[m_{1}]\rangle\otimes\dots\otimes|\Phi[m_{n}]\rangle$ which represents the state of $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$. This state indicates that the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ record the values $m_{1}\dots m_{n}$ as the measurement outcomes over $\mathcal{S}_{1},\dots,\mathcal{S}_{n}$, respectively.
In the superposition (6), on letting $n\to\infty$, the length of the records $m_{1}\dots m_{n}$ of the values as the measurement outcomes in the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ diverges to infinity. The consideration of this limiting case results in the definition of a world. Namely, a world is defined as an infinite sequence of records of the values as the measurement outcomes in the apparatuses. Thus, in the case described so far, a world is an infinite sequence over $\Omega$, and the finite records $m_{1}\dots m_{n}$ in each state $|\Phi[m_{1}\dots m_{n}]\rangle$ in the superposition (6) of the total system is a prefix of a world.
Then, for aiming at deriving Postulate 4, MWI assigns “weight” to each of worlds. Namely, it introduces a probability measure on the set of all worlds in the following manner: First, MWI introduces a probability measure representation on the set of prefixes of worlds, i.e., the set $\Omega^{*}$ in this case. This probability measure representation is given by a function $r\colon\Omega^{*}\to[0,1]$ with
$r(m_{1}\dotsc m_{n})=\prod_{k=1}^{n}\langle\Psi_{k}|M_{m_{k}}^{\dagger}M_{m_{k}}|\Psi_{k}\rangle,$ (7)
which is the square of the norm of each state $(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}\dots m_{n}]\rangle$ in the superposition (6). Using the completeness equation (4), it is easy to check that $r$ is certainly a probability measure representation over $\Omega$. We call the probability measure representation $r$ the probability measure representation for the prefixes of worlds. Then MWI tries to derive Postulate 4 by adopting the probability measure induced by the probability measure representation $r$ for the prefixes of worlds as the probability measure on the set of all worlds.
We summarize the above consideration and clarify the definitions of the notion of world and the notion of the probability measure representation for the prefixes of worlds as in the following.
Definition 6.1 (The probability measure representation for the prefixes of worlds, Tadaki [47]).
Consider an arbitrary finite-dimensional quantum system $\mathcal{S}$ and a measurement over $\mathcal{S}$ described by arbitrary measurement operators $\{M_{m}\}_{m\in\Omega}$, where the measurement process is described by (5) as an interaction of the system $\mathcal{S}$ with an apparatus $\mathcal{A}$. We suppose the following situation:
  1. (i)
    There are countably infinite copies $\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\dotsc$ of the system $\mathcal{S}$ and countably infinite copies $\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3},\dotsc$ of the apparatus $\mathcal{A}$.
  2. (ii)
    For each $n\in\mathbb{N}^{+}$, the system $\mathcal{S}_{n}$ is prepared in a state $|\Psi_{n}\rangle$, [^2] In Definition 6.1, all $|\Psi_{n}\rangle$ are not required to be an identical state. while the apparatus $\mathcal{A}_{n}$ is prepared in a state $|\Phi_{\mathrm{init}}\rangle$, and then the measurement described by $\{M_{m}\}_{m\in\Omega}$ is performed over $\mathcal{S}_{n}$ by interacting it with the apparatus $\mathcal{A}_{n}$ according to the unitary time-evolution (5).
  3. (iii)
    Starting the measurement described by $\{M_{m}\}_{m\in\Omega}$ over $\mathcal{S}_{1}$, the measurement described by $\{M_{m}\}_{m\in\Omega}$ over each $\mathcal{S}_{n}$ is performed in the increasing order of $n$.
We then note that, for each $n\in\mathbb{N}^{+}$, immediately after the measurement described by $\{M_{m}\}_{m\in\Omega}$ over $\mathcal{S}_{n}$, the state of the total system consisting of $\mathcal{S}_{1},\mathcal{A}_{1},\mathcal{S}_{2},\mathcal{A}_{2},\dots,\mathcal{S}_{n},\mathcal{A}_{n}$ is
$|\Theta_{n}\rangle:=\sum_{m_{1},\dots,m_{n}\in\Omega}|\Theta(m_{1},\dots,m_{n})\rangle,$
where
$|\Theta(m_{1},\dots,m_{n})\rangle:=(M_{m_{1}}|\Psi_{1}\rangle)\otimes\dots\otimes(M_{m_{n}}|\Psi_{n}\rangle)\otimes|\Phi[m_{1}]\rangle\otimes\dots\otimes|\Phi[m_{n}]\rangle.$
The vectors $M_{m_{1}}|\Psi_{1}\rangle,\dots,M_{m_{n}}|\Psi_{n}\rangle$ are states of $\mathcal{S}_{1},\dots,\mathcal{S}_{n}$, respectively, and the vectors
$|\Phi[m_{1}]\rangle,\dots,|\Phi[m_{n}]\rangle$
are states of $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, respectively. The state vector $|\Theta_{n}\rangle$ of the total system is normalized while each of the vectors $\{|\Theta(m_{1},\dots,m_{n})\rangle\}_{m_{1},\dots,m_{n}\in\Omega}$ is not necessarily normalized. Then, the probability measure representation for the prefixes of worlds is defined as a function $p\colon\Omega^{*}\to[0,1]$ such that
$p(m_{1}\dotsc m_{n})=\langle\Theta(m_{1},\dots,m_{n})|\Theta(m_{1},\dots,m_{n})\rangle$ (8)
for every $n\in\mathbb{N}^{+}$ and every $m_{1},\dots,m_{n}\in\Omega$. Moreover, an infinite sequence over $\Omega$, i.e., an infinite sequence of possible outcomes of the measurement described by $\{M_{m}\}_{m\in\Omega}$, is called a world. ∎
In Definition 6.1, it is easy to check that the function $p$ defined by (8) is certainly a probability measure representation over $\Omega$.
As mentioned above, the original MWI by Everett [22] aimed to derive Postulate 4 (i), the Born rule, from the remaining postulates. However, it would seem impossible to do this for several reasons (see Tadaki [47] for these reasons). Instead, it is appropriate to introduce an additional postulate in the rigorous framework of MWI developed above, in order to overcome the defect of the original MWI and to make quantum mechanics operationally perfect. Thus, we put forward the principle of typicality as follows.
Intuitively, the principle of typicality states that our world is typical. Namely, our world is Martin-Löf random with respect to the probability measure on the set of all worlds, induced by the probability measure representation for the prefixes of worlds, in the superposition of the total system which consists of systems being measured and apparatuses measuring them. Formally, the principle of typicality is stated as the following postulate.
Postulate 5 (The principle of typicality, Tadaki [42]).
We assume the notation and notions introduced and defined in Definition 6.1, and consider the same setting of measurements as considered in Definition 6.1. Let $\omega\in\Omega^{\infty}$ be our world. Then $\omega$ is Martin-Löf random with respect to the probability measure $\mu_{p}$ induced by the probability measure representation $p\colon\Omega^{*}\to[0,1]$ for the prefixes of worlds. Moreover, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}_{n}+\mathcal{A}_{n}$, which consists of the $n$th copy $\mathcal{S}_{n}$ of the system $\mathcal{S}$ and the $n$th copy $\mathcal{A}_{n}$ of the apparatus $\mathcal{A}$, immediately after the measurement by $\mathcal{A}_{n}$ performed over $\mathcal{S}_{n}$ is given by
$(M_{\omega(n)}|\Psi_{n}\rangle)\otimes|\Phi[\omega(n)]\rangle,$
up to the normalization factor. ∎
For the comprehensive arguments of the validity of Postulate 5, the principle of typicality, see Tadaki [47]. For example, based on the results of Tadaki [37, 38, 44], we can see that Postulate 5 is certainly a refinement of Postulate 4 from the point of view of our intuitive understanding of the notion of probability.
7 Apparatuses measuring apparatuses and their confirming points
In this section we further develop the rigorous framework based on the principle of typicality reviewed in the preceding section, in order to make it applicable to the situation where apparatuses perform measurements over other apparatuses as well as over a normal quantum system being measured.
7.1 The general setting of the problem
Consider measuring apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ performing measurements $\mathcal{M}^{1},\dots,\mathcal{M}^{n}$ described by measurement operators $\{M^{1}_{m_{1}}\}_{m_{1}\in\Omega_{1}},\dots,\{M^{n}_{m_{n}}\}_{m_{n}\in\Omega_{n}}$, respectively, where $\Omega_{1},\dots,\Omega_{n}$ are the sets of all possible outcomes of the measurements $\mathcal{M}^{1},\dots,\mathcal{M}^{n}$. The measuring apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ themselves are quantum systems with state spaces $\overline{\mathcal{H}}_{1}\dots,\overline{\mathcal{H}}_{n}$, respectively. We also consider a quantum system $\mathcal{S}$ with state space $\mathcal{H}_{\mathcal{S}}$. We assume the situation that, for each $k=1,\dots,n$, the apparatus $\mathcal{A}_{k}$ performs the measurement $\mathcal{M}^{k}$ over the composite system consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{k-1}$. Therefore, we assume that, for every $k=1,\dots,n$ and every $m_{k}\in\Omega_{k}$, the measurement operator $M^{k}_{m_{k}}$ is an endomorphism of $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{k-1}$.
Thus, according to (5), for each $k=1,\dots,n$, the measurement process of the measurement $\mathcal{M}^{k}$ is described by a unitary operator $U_{k}$ acting on $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{k}$ such that
$U_{k}(|\Psi\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle)=\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Psi\rangle)\otimes|\Phi^{k}[m_{k}]\rangle$ (9)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{k-1}$. For each $k=1,\dots,n$, the vector $|\Phi^{k}_{\mathrm{init}}\rangle\in\overline{\mathcal{H}}_{k}$ is the initial state of the apparatus $\mathcal{A}_{k}$, and the vector $|\Phi^{k}[m_{k}]\rangle\in\overline{\mathcal{H}}_{k}$ is a final state of the apparatus $\mathcal{A}_{k}$ for each $m_{k}\in\Omega_{k}$, with $\langle\Phi^{k}[m_{k}]|\Phi^{k}[m_{k}^{\prime}]\rangle=\delta_{m_{k},m_{k}^{\prime}}$. For every $k=1,\dots,n$ and every $m_{k}\in\Omega_{k}$, the state $|\Phi^{k}[m_{k}]\rangle$ indicates that the apparatus $\mathcal{A}_{k}$ records the value $m_{k}$ as the measurement outcome. For each $k=1,\dots,n$, let $\mathcal{H}_{k}$ be a subspace of $\overline{\mathcal{H}}_{k}$ spanned by the vectors
$\{|\Phi^{k}[m_{k}]\rangle\}_{m_{k}\in\Omega_{k}}.$
As a natural requirement, for each $k=1,\dots,n$, we assume that
$M^{k}_{m_{k}}|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$ (10)
for every $m_{k}\in\Omega_{k}$ and every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{k}\otimes\dots\otimes\mathcal{H}_{k-1}$.
Let us consider the total system $\mathcal{S}_{\mathrm{Total}}$ consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$. For any $i,j\in\{0,\dots,n\}$ with $i\leq j$, we use $I_{i,\dots,j}$ to denote the identity operator acting on $\overline{\mathcal{H}}_{i}\otimes\dots\otimes\overline{\mathcal{H}}_{j}$, where $\overline{\mathcal{H}}_{0}$ denotes $\mathcal{H}_{\mathcal{S}}$. Therefore $I_{0}$ denotes the identity operator acting on $\mathcal{H}_{\mathcal{S}}$, for example. For each $k=1,\dots,n$, we define a unitary operator $U_{\mathrm{Total}}^{1,\dots,k}$ on $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{n}$ by
$U_{\mathrm{Total}}^{1,\dots,k}:=(U_{k}\otimes I_{k+1,\dots,n})\circ(U_{k-1}\otimes I_{k,\dots,n})\circ\dotsm\dotsm\circ(U_{2}\otimes I_{3,\dots,n})\circ(U_{1}\otimes I_{2,\dots,n}).$
Since $U_{1},\dots,U_{n}$ are unitary operators, based on (9) and (10), it is easy to see that for every $k=1,\dots,n$ there exist unique measurement operators $\{M^{1,\dots,k}_{m_{1},\dots,m_{k}}\}_{(m_{1},\dots,m_{k})\in\Omega_{1}\times\dots\times\Omega_{k}}$, satisfying the completeness equation
$\sum_{(m_{1},\dots,m_{k})\in\Omega_{1}\times\dots\times\Omega_{k}}{M^{1,\dots,k}_{m_{1},\dots,m_{k}}}^{\dagger}M^{1,\dots,k}_{m_{1},\dots,m_{k}}=I_{0},$
such that for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ it holds that
$\begin{split}&U_{\mathrm{Total}}^{1,\dots,k}(|\Psi\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle)\\ &=\sum_{(m_{1},\dots,m_{k})\in\Omega_{1}\times\dots\times\Omega_{k}}(M^{1,\dots,k}_{m_{1},\dots,m_{k}}|\Psi\rangle)\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.\end{split}$ (11)
For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if $|\Psi\rangle$ is a state of the system $\mathcal{S}$ immediately before the measurement $\mathcal{M}^{1}$, then $|\Psi\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$ is the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately before the measurement $\mathcal{M}^{1}$ and $U_{\mathrm{Total}}^{1,\dots,k}(|\Psi\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle)$ is the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{k}$ for every $k=1,\dots,n$.
Let $|\Psi^{0}_{\mathrm{init}}\rangle$ be an arbitrary state of the system $\mathcal{S}$ immediately before the measurement $\mathcal{M}^{1}$. Then, for each $k=1,\dots,n$, we use $|\Psi^{k}_{\mathrm{Total}}\rangle$ to denote the state $U_{\mathrm{Total}}^{1,\dots,k}(|\Psi^{0}_{\mathrm{init}}\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle)$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{k}$. In particular, it follows from (11) that the state $|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n}$ is given by
$\begin{split}|\Psi^{n}_{\mathrm{Total}}\rangle=\sum_{(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}}(M^{1,\dots,n}_{m_{1},\dots,m_{n}}|\Psi^{0}_{\mathrm{init}}\rangle)\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle.\end{split}$ (12)
In our framework based on the principle of typicality, the unitary operator $U_{\mathrm{Total}}^{1,\dots,n}$ applying to the initial state $|\Psi^{0}_{\mathrm{init}}\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurements $\mathcal{M}^{1},\dots,\mathcal{M}^{n}$ in this order is infinitely repeated. It follows from (12) that the application of $U_{\mathrm{Total}}^{1,\dots,n}$ can be regarded as a single measurement which is described by the measurement operators $\{M^{1,\dots,n}_{m_{1},\dots,m_{n}}\}_{(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}}$ and whose all possible outcomes form the set $\Omega_{1}\times\dots\times\Omega_{n}$. Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Thus, according to Definition 6.1, we can see that a world is an infinite sequence over $\Omega_{1}\times\dots\times\Omega_{n}$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{P}$ on $(\Omega_{1}\times\dots\times\Omega_{n})^{\infty}$, where $P$ is a finite probability space on $\Omega_{1}\times\dots\times\Omega_{n}$ such that $P(m_{1},\dots,m_{n})$ is the square of the norm of the vector
$(M^{1,\dots,n}_{m_{1},\dots,m_{n}}|\Psi^{0}_{\mathrm{init}}\rangle)\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle$
for every $(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}$. Here $\Omega_{1}\times\dots\times\Omega_{n}$ is the set of all possible records of the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ in the repeated once of the experiments.
Thus, we can apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurement described by the measurement operators $\{M^{1,\dots,n}_{m_{1},\dots,m_{n}}\}_{(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}}$ in the above setting. This $\omega$ is an infinite sequence over $\Omega_{1}\times\dots\times\Omega_{n}$ consisting of records in the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ which is being generated by the infinite repetition of the measurement described by the measurement operators $\{M^{1,\dots,n}_{m_{1},\dots,m_{n}}\}_{(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $(\Omega_{1}\times\dots\times\Omega_{n})^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\omega^{1},\dots,\omega^{n}$ to denote the infinite sequences over $\Omega_{1},\dots,\Omega_{n}$, respectively, such that $(\omega^{1}(l),\dots,\omega^{n}(l))=\omega(l)$ for every $l\in\mathbb{N}^{+}$. Then, it follows from Postulate 5 and (12) that, in our world $\omega$, for each $l\in\mathbb{N}^{+}$ the state of the $l$th total system $\mathcal{S}_{\mathrm{Total}}$, i.e., the $l$th copy of a composite system consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n}$ is given by
$(M^{1,\dots,n}_{\omega^{1}(l),\dots,\omega^{n}(l)}|\Psi^{0}_{\mathrm{init}}\rangle)\otimes|\Phi^{1}[\omega^{1}(l)]\rangle\otimes\dots\otimes|\Phi^{n}[\omega^{n}(l)]\rangle,$ (13)
up to the normalization factor.
Now, originally, the state $|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ is a superposition of the vectors of the form
$(M^{1,\dots,n}_{m_{1},\dots,m_{n}}|\Psi^{0}_{\mathrm{init}}\rangle)\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle$ (14)
immediately after the measurement $\mathcal{M}^{n}$ according to (12). However, as we saw in the above argument leading to the state (13), the application of the principle of typicality implies that the state of the total system $\mathcal{S}_{\mathrm{Total}}$ is actually given by a specific vector of the form of (14) (up to the normalization factor) immediately after the measurement $\mathcal{M}^{n}$ in each repetition of the measurement described by the measurement operators $\{M^{1,\dots,n}_{m_{1},\dots,m_{n}}\}_{(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}}$.
In what follows, we study the problem to determine the whole state or the respective states of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{k}$ in the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after each measurement $\mathcal{M}^{k}$ with $1\leq k<n$, provided that the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n}$ is specified as a vector of the form
$|\Psi\rangle\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle$
with $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ and $(m_{1},\dots,m_{n})\in\Omega_{1}\times\dots\times\Omega_{n}$, just like the vector (13). In order to solve this problem, we must introduce some postulates. For that purpose, we first introduce some terminology.
7.2 Confirming points for the final states of apparatuses
Definition 7.1.
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{j}\in\Omega_{j}$, every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exists a vector $|\Theta_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
$M^{j}_{m_{j}}(|\Psi\rangle\otimes|\Psi_{R}\rangle)=|\Psi\rangle\otimes|\Theta_{R}\rangle.$
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the states of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
Definition 7.2 (Confirming point for the states of system).
The notion of the confirming point for the states of the system $\mathcal{S}$ is defined as follows: Let $V$ be the set of all $j\in\{1,\dots,n\}$ such that the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V\neq\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the system $\mathcal{S}$ itself if $j=1$ and the apparatus $\mathcal{A}_{j-1}$ otherwise, where $j:=\min V$. If $V=\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the apparatus $\mathcal{A}_{n}$. ∎
Definition 7.3.
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{i}\in\Omega_{i}$, every $m_{j}\in\Omega_{j}$, every $|\Psi_{L}\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exist a set $\{|\Theta_{L}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Theta_{R}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
$M^{j}_{m_{j}}(|\Psi_{L}\rangle\otimes|\Phi^{i}[m_{i}]\rangle\otimes|\Psi_{R}\rangle)=\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m_{i}]\rangle\otimes|\Theta_{R}[l]\rangle.$
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
Definition 7.4 (Confirming point for the final states of apparatus).
Let $i\in\{1,\dots,n\}$. The notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as follows: Let $V_{i}$ be the set of all $j\in\{i+1,\dots,n\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V_{i}\neq\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{j-1}$ with $j:=\min V_{i}$. If $V_{i}=\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{n}$. Thus, the confirming point for the final states of the apparatus $\mathcal{A}_{n}$ is the apparatus $\mathcal{A}_{n}$ itself. ∎
Consider the case where there are only two apparatuses $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with $n=2$ and the final states of the apparatus $\mathcal{A}_{1}$ is not unchanged after the measurement by the apparatus $\mathcal{A}_{2}$. In this case, we cannot think the measurement $\mathcal{M}^{1}$ by the apparatus $\mathcal{A}_{1}$ to be completed immediately after the measurement $\mathcal{M}^{1}$ itself, since the state of the apparatus $\mathcal{A}_{1}$ is disturbed by the subsequent measurement $\mathcal{M}^{2}$ by the apparatus $\mathcal{A}_{2}$. However, it follows from (11) that
$\begin{split}U_{\mathrm{Total}}^{1,2}(|\Psi\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes|\Phi^{2}_{\mathrm{init}}\rangle)=\sum_{(m_{1},m_{2})\in\Omega_{1}\times\Omega_{2}}(M^{1,2}_{m_{1},m_{2}}|\Psi\rangle)\otimes|\Phi^{1}[m_{1}]\rangle\otimes|\Phi^{2}[m_{2}]\rangle\end{split}$
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. This implies that the combination of the apparatus $\mathcal{A}_{1}$ with the apparatus $\mathcal{A}_{2}$ can be regarded as a single measurement, which is described by the measurement operators $\{M^{1,2}_{m_{1},m_{2}}\}_{(m_{1},m_{2})\in\Omega_{1}\times\Omega_{2}}$ and whose all possible outcomes form the set $\Omega_{1}\times\Omega_{2}$. Thus, we can say that the measurement $\mathcal{M}^{1}$ by the apparatus $\mathcal{A}_{1}$ is really completed at the instant when the measurement $\mathcal{M}^{2}$ by the apparatus $\mathcal{A}_{2}$ is completed. The generalization of this idea to the general case where there are $n$ apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ with an arbitrary positive integer $n$ is the notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ with an arbitrary integer $i\in\{1,\dots,n\}$, introduced in Definition 7.4 above. We think that the measurement by each apparatus $\mathcal{A}_{i}$ is really completed at the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ itself.
Now, to motivate Postulate 6 below, we first prove the following proposition.
Proposition 7.5.
Without assuming Postulate 5, based on (9) and (10) the following (i) and (ii) hold:
  1. (i)
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
  2. (ii)
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$.
Proof.
(i) Let $i\in\{1,\dots,n\}$, and let $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. Suppose that the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$. Then, in order to prove Proposition 7.5 (i), it is sufficient to show that, for every $k=i,\dots,n$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
Now, for an arbitrary $k=i,\dots,n$, assume that the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$. It follows from (11) that the state $|\Psi^{k-1}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately before the measurement $\mathcal{M}^{k}$ has the form
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=|\Psi^{k-1}\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$
for some state $|\Psi^{k-1}\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$. Let $\{|\Psi[l]\rangle\}_{l}$ be an orthonormal basis of $\mathcal{H}_{\mathcal{S}}$ such that $|\Psi[1]\rangle=|\Psi\rangle$. Then the vector $|\Psi^{k-1}\rangle$ can be written in the following form:
$|\Psi^{k-1}\rangle=\sum_{l}|\Psi[l]\rangle\otimes|\Theta_{R}[l]\rangle,$ (15)
where $\{|\Theta_{R}[l]\rangle\}_{l}$ is a set of vectors in $\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$. For each $l$, we use $|\Gamma[l]\rangle$ to denote the vector
$|\Psi[l]\rangle\otimes|\Theta_{R}[l]\rangle.$
Since the state of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exists a set $\{|\Upsilon_{R}[l,m_{k}]\rangle\}_{l,m_{k}}$ of vectors in $\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $l$ and every $m_{k}\in\Omega_{k}$,
$M^{k}_{m_{k}}|\Gamma[l]\rangle=|\Psi[l]\rangle\otimes|\Upsilon_{R}[l,m_{k}]\rangle.$ (16)
Thus it follows from (9) that the state $|\Psi^{k}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{k}$ has the following form:
$\begin{split}|\Psi^{k}_{\mathrm{Total}}\rangle&=(U_{k}(|\Psi^{k-1}\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle))\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Psi^{k-1}\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{l}\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Gamma[l]\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\end{split}$ (17)
Since the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (16) that $M^{k}_{m_{k}}|\Gamma[l]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $l\geq 2$. Thus, using the completeness relation
$\sum_{m_{k}\in\Omega_{k}}{M^{k}_{m_{k}}}^{\dagger}M^{k}_{m_{k}}=I_{0,1,\dots,k-1},$
we have that $|\Gamma[l]\rangle=0$ for every $l\geq 2$. It follows from (15) that the state $|\Psi^{k-1}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately before the measurement $\mathcal{M}^{k}$ has the form
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=|\Psi\rangle\otimes|\Theta_{R}[1]\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.$
Hence, the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{k}$, as desired.
(ii) Let $i,j\in\{1,\dots,n\}$ with $i<j$, and let $m_{i}\in\Omega_{i}$. Suppose that the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. Then, in order to prove Proposition 7.5 (ii), it is sufficient to show that, for every $k=j,\dots,n$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
Now, for an arbitrary $k=j,\dots,n$, assume that the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$. It follows from (11) that the state $|\Psi^{k-1}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately before the measurement $\mathcal{M}^{k}$ has the form
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=|\Psi\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$
for some state $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$. Then the vector $|\Psi\rangle$ can be written in the following form:
$|\Psi\rangle=\sum_{m\in\Omega_{i}}\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m]\rangle\otimes|\Theta_{R}[m,l]\rangle,$ (18)
where $\{|\Theta_{L}[l]\rangle\}_{l}$ is a set of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and $\{|\Theta_{R}[m,l]\rangle\}_{m,l}$ is a set of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{k-1}$. For each $m\in\Omega_{i}$, we use $|\Gamma[m]\rangle$ to denote the vector
$\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m]\rangle\otimes|\Theta_{R}[m,l]\rangle.$
Since the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exist a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $m\in\Omega_{i}$ and every $m_{k}\in\Omega_{k}$,
$M^{k}_{m_{k}}|\Gamma[m]\rangle=\sum_{h}|\Upsilon_{L}[m,m_{k},h]\rangle\otimes|\Phi^{i}[m]\rangle\otimes|\Upsilon_{R}[m,m_{k},h]\rangle.$ (19)
Thus it follows from (9) that the state $|\Psi^{k}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{k}$ has the following form:
$\begin{split}|\Psi^{k}_{\mathrm{Total}}\rangle&=(U_{k}(|\Psi\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle))\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Psi\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m\in\Omega_{i}}\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Gamma[m]\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\end{split}$ (20)
Since the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (19) that $M^{k}_{m_{k}}|\Gamma[m]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $m\in\Omega_{i}$ with $m\neq m_{i}$. Thus, using the completeness relation
$\sum_{m_{k}\in\Omega_{k}}{M^{k}_{m_{k}}}^{\dagger}M^{k}_{m_{k}}=I_{0,1,\dots,k-1},$
we have that $|\Gamma[m]\rangle=0$ for every $m\in\Omega_{i}$ with $m\neq m_{i}$. It follows from (18) that the state $|\Psi^{k-1}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately before the measurement $\mathcal{M}^{k}$ has the form
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m_{i}]\rangle\otimes|\Theta_{R}[m,l]\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.$
Hence, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{k}$, as desired. ∎
7.3 Postulates for determining the states of apparatuses in early stages
In this section, we have so far made an argument based on Postulate 3, the unitary time-evolution, and not based on Postulate 5, the principle of typicality. The application of Postulate 5 results in the non-unitary time-evolution in the measurement process, as we saw in Section 6. In the rest of this paper, we regard the states $|\Psi^{1}_{\mathrm{Total}}\rangle,|\Psi^{2}_{\mathrm{Total}}\rangle,\dots,|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurements $\mathcal{M}^{1},\mathcal{M}^{2},\dots,\mathcal{M}^{n}$, respectively, as virtual states which provide a useful means for the calculation in our framework based on the principle of typicality, whereas the initial state $|\Psi^{0}_{\mathrm{init}}\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ is treated as a real state. Actually, the state $|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the whole measurement has been used as a virtual state in our framework based on the principle of typicality, developed in Section 6. However, suggested by Proposition 7.5, we propose the following postulate for real states.
Postulate 6.
  1. (i)
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
  2. (ii)
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$. ∎
We remark the following: The statement of Postulate 6 appears to be the same as that of Proposition 7.5 on the surface. However, Proposition 7.5 assumes the unitary time-evolution in the total measurement process consisting of the measurements $\mathcal{M}^{1},\mathcal{M}^{2},\dots,\mathcal{M}^{n}$, as we saw above. In contrast, Postulate 6 does not assume the unitary time-evolution in the total measurement process. Postulate 6 plays a major role in the analysis of Wigner’s friend in the subsequent sections.
Next, we consider the problem to determine the whole state of the total system $\mathcal{S}_{\mathrm{Total}}$, consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n-1}$, given the whole state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n}$.
Let $k\in\mathbb{N}$, and let $i_{1},\dots,i_{k}$ be integers such that $1\leq i_{1}<\dots<i_{k}\leq n$. Let $|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle$ be states of the apparatuses $\mathcal{A}_{i_{1}},\dots,\mathcal{A}_{i_{k}}$, respectively. In the case of $k\geq 1$, we define
$P(|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle)$
as a projector onto the subspace of the state space of the total system $\mathcal{S}_{\mathrm{Total}}$ where the states of the apparatuses $\mathcal{A}_{i_{1}},\dots,\mathcal{A}_{i_{k}}$ are determined as $|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle$, respectively. Formally, in the case of $k\geq 1$, $P(|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle)$ is defined as a projector acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{n}$ of the total system $\mathcal{S}_{\mathrm{Total}}$ of the form
$I_{0,1,\dots,i_{1}-1}$ $\otimes\lvert\Phi^{i_{1}}\rangle\langle\Phi^{i_{1}}\rvert\otimes I_{i_{1}+1,\dots,i_{2}-1}\otimes\lvert\Phi^{i_{2}}\rangle\langle\Phi^{i_{2}}\rvert\otimes I_{i_{2}+1,\dots,i_{3}-1}\otimes\dotsm\dotsm$
$\dotsm\dotsm\otimes I_{i_{k-2}+1,\dots,i_{k-1}-1}\otimes\lvert\Phi^{i_{k-1}}\rangle\langle\Phi^{i_{k-1}}\rvert\otimes I_{i_{k-1}+1,\dots,i_{k}-1}\otimes\lvert\Phi^{i_{k}}\rangle\langle\Phi^{i_{k}}\rvert\otimes I_{i_{k}+1,\dots,n}.$
In the case of $k=0$, we define $P(|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle)$, i.e., $P()$, as the identity operator $I_{0,1,\dots,n}$ acting on the whole state space $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{n}$ of the total system $\mathcal{S}_{\mathrm{Total}}$.
Postulate 7.
Suppose that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. Suppose that the state of the total system $\mathcal{S}_{\mathrm{Total}}$, consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n}$ is
$|\Psi\rangle\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle,$
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Let $A$ be the set of $i\in\{1,\dots,n-1\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{n}$. We define a finite sequence $i_{1},i_{2},\dots,i_{k}\in\{1,2,\dots,n-1\}$ by the condition that $A=\{i_{1},i_{2},\dots,i_{k}\}$ and $i_{1}<i_{2}<\dots<i_{k}$. Then the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
$P(|\Phi^{i_{1}}[m_{i_{1}}]\rangle,\dots,|\Phi^{i_{k}}[m_{i_{k}}]\rangle)|\Psi^{n-1}_{\mathrm{Total}}\rangle,$ (21)
up to the normalization factor. Note that this state is just $|\Psi^{n-1}_{\mathrm{Total}}\rangle$ in the case of $A=\emptyset$, i.e., in the case of $k=0$. ∎
Postulate 7 only states the relation between the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n}$ and the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$. However, as we will see in Section 14.1, we can use Postulate 7 in a recursive manner to state the relation between the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n}$ and the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{i}$ with $1\leq i<n$.
In the rest of this section, let us investigate immediate consequences of Postulate 6 and Postulate 7, and thus investigate the validity of these postulates. For that purpose, we consider the normal case where, for each $i=1,\dots,n-1$, the apparatus $\mathcal{A}_{i}$ performs the measurement $\mathcal{M}^{i}$ only over the system $\mathcal{S}$ and not over any of the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{i-1}$. Thus, we assume that, for every $i=2,\dots,n$ and every $m_{i}\in\Omega_{i}$, the measurement operator $M^{i}_{m_{i}}$ has the following form:
$M^{i}_{m_{i}}=L^{i}_{m_{i}}\otimes I_{1,\dots,i-1},$
where $\{L^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ are measurement operators acting on $\mathcal{H}_{\mathcal{S}}$. It is then easy to see that, for every $i=1,\dots,n-1$, the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i+1}$. Thus, for every $i=1,\dots,n-1$, the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is the apparatus $\mathcal{A}_{i}$ itself.
Suppose that the state of the total system $\mathcal{S}_{\mathrm{Total}}$, consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n}$ is
$|\Psi\rangle\otimes|\Phi^{1}[m_{1}]\rangle\otimes\dots\otimes|\Phi^{n}[m_{n}]\rangle,$
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Then, according to Postulate 6 we have that, for every $i=1,\dots,n-1$, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{i+1}$, that is, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{i}$. This result is comprehensively used in our former works implicitly to make the operational refinements of quantum information theory in our framework based on the principle of typicality [46, 47, 49, 50].
Suppose further that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. Then, since the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{n}$ for every $i=1,\dots,n-1$, according to Postulate 7 we have that the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
$P(|\Phi^{1}[m_{1}]\rangle,\dots,|\Phi^{n-1}[m_{n-1}]\rangle)|\Psi^{n-1}_{\mathrm{Total}}\rangle,$ (22)
up to the normalization factor. This result is also comprehensively used in our former works implicitly to make the operational refinements of quantum information theory in our framework based on the principle of typicality [46, 47, 49, 50]. In particular, the result (22) implicitly played a crucial role in the derivations of Postulate 8 of Tadaki [47], which is a refined rule of the Born rule for mixed states, from the principle of typicality in several scenarios of the setting of measurements, as described in Sections 11 and 12 of Tadaki [47]. Thus, this effectiveness of the result (22) supports general use of Postulate 7 in a specific situation such as Wigner’s friend and Deutsch’s thought experiment.
The reason why the measurement operators $\{M^{1}_{m_{1}}\}_{m_{1}\in\Omega_{1}},\dots,\{M^{n}_{m_{n}}\}_{m_{n}\in\Omega_{n}}$ must be restricted to PVMs in Postulate 7 is the following: If we allow $\{M^{1}_{m_{1}}\}_{m_{1}\in\Omega_{1}},\dots,\{M^{n}_{m_{n}}\}_{m_{n}\in\Omega_{n}}$ to be general measurement operators then Postulate 7 becomes too strong and results in a contradiction. We present the details of this argument in Remark 14.2 in Section 14.1. For another justification of Postulate 7, see Remark 14.1.
In the subsequent sections, we apply Postulates 6 and 7 to several examples of a general case where the apparatus $\mathcal{A}_{i}$ performs the measurement $\mathcal{M}^{i}$ over the composite system consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{i-1}$ for each $i=1,\dots,n$.
8 Wigner’s friend
In this section, we review the Wigner’s friend paradox [54] in the terminology of the conventional quantum mechanics.
Consider a single qubit system $\mathcal{S}$ with state space $\mathcal{H}_{\mathcal{S}}$. An observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ described by the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ over the system $\mathcal{S}$ (according to Postulate 4), where
$M^{\mathcal{F}}_{0}:=\lvert 0\rangle\langle 0\rvert\quad\text{and}\quad M^{\mathcal{F}}_{1}:=\lvert 1\rangle\langle 1\rvert$, (23)
and $|0\rangle$ and $|1\rangle$ form an orthonormal basis of the state space $\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$.
Hereafter, we assume that the observer $\mathcal{F}$ itself is a quantum system with state space $\mathcal{H}_{\mathcal{F}}$. Then we consider an observer $\mathcal{W}$ who is on the outside of the composite system $\mathcal{S}+\mathcal{F}$ consisting of the system $\mathcal{S}$ and the observer $\mathcal{F}$. The observer $\mathcal{W}$ can also be regarded as an environment for the composite system $\mathcal{S}+\mathcal{F}$. Then, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is described by a unitary operator $U_{\mathrm{I}}$ such that
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (24)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. The unitary operator $U_{\mathrm{I}}$ describes the interaction between the system $\mathcal{S}$ and the observer $\mathcal{F}$ as a quantum system. The vector $|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{F}}$ is the initial state of the observer $\mathcal{F}$, and $|\Phi^{\mathcal{F}}[k]\rangle\in\mathcal{H}_{\mathcal{F}}$ is a final state of the observer $\mathcal{F}$ for each $k=0,1$, with $\langle\Phi^{\mathcal{F}}[k]|\Phi^{\mathcal{F}}[k^{\prime}]\rangle=\delta_{k,k^{\prime}}$. For each $k=0,1$, the state $|\Phi^{\mathcal{F}}[k]\rangle$ indicates that the observer $\mathcal{F}$ records the value $k$. The unitary time-evolution (24) is the quantum mechanical description of the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ over the composite system $\mathcal{S}+\mathcal{F}$, from the point of view of the observer $\mathcal{W}$.
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
$\begin{split}M^{\mathcal{W}}_{0}&:=\lvert\Phi^{\mathcal{F}}[0]\rangle\langle\Phi^{\mathcal{F}}[0]\rvert,\\ M^{\mathcal{W}}_{1}&:=\lvert\Phi^{\mathcal{F}}[1]\rangle\langle\Phi^{\mathcal{F}}[1]\rvert,\\ M^{\mathcal{W}}_{2}&:=\sqrt{I_{\mathcal{F}}-{M^{\mathcal{W}}_{0}}^{\dagger}M^{\mathcal{W}}_{0}-{M^{\mathcal{W}}_{1}}^{\dagger}M^{\mathcal{W}}_{1}}=I_{\mathcal{F}}-M^{\mathcal{W}}_{0}-M^{\mathcal{W}}_{1},\end{split}$
and $I_{\mathcal{F}}$ denotes the identity operator acting on $\mathcal{H}_{\mathcal{F}}$.
Now, let us assume that the system $\mathcal{S}$ is initially in the state $|+\rangle$ in the above setting, where
$|+\rangle:=\frac{1}{\sqrt{2}}\left(|0\rangle+|1\rangle\right).$ (25)
Then, we have the following two cases, depending on how we treat the measurement $\mathcal{M}^{\mathcal{F}}$:
Case 1.
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
  1. (i)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|0\rangle$ and $|\Phi^{\mathcal{F}}[0]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle$.
  2. (ii)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|1\rangle$ and $|\Phi^{\mathcal{F}}[1]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle$.
Here, each possibility occurs with the same probability $1/2$.
Case 2.
According to (24), immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ consisting of $\mathcal{S}$ and $\mathcal{F}$ is
$\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$ (26)
Thus, the state (26) of the composite system $\mathcal{S}+\mathcal{F}$ in Case 2 is different from neither $|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle$ in the possibility (i) nor $|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle$ in the possibility (ii) in Case 1, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$. This inconsistency is the content of the paradox in the Wigner’s friend paradox [54]. Note that, even if the composite system $\mathcal{S}+\mathcal{F}$ is in the state (26) immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, according to Postulate 4 the subsequent measurement $\mathcal{M}^{\mathcal{W}}$ results in one of the two possibilities (i) and (ii) above with the same probability $1/2$ for the post-measurement state of the composite system $\mathcal{S}+\mathcal{F}$, just as in Case 1.
9 Analysis of Wigner’s friend based on the principle of typicality
In this section, we make an analysis of the Wigner’s friend paradox, described in the preceding section, in our framework of quantum mechanics based on Postulates 5, 6, and 7, together with Postulates 1, 2, and 3. We will then see that the measurement $\mathcal{M}^{\mathcal{F}}$ settles the composite system $\mathcal{S}+\mathcal{F}$ down into one of the two possibilities (i) and (ii) in Case 1 above, and not into the ‘undecided’ state (26) in Case 2.
To proceed this program, first of all we have to implement everything, i.e., the two measurements $\mathcal{M}^{\mathcal{F}}$ and $\mathcal{M}^{\mathcal{W}}$ in the setting of the Wigner’s friend paradox, by unitary time-evolution. The measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is already described by the unitary operator $U_{\mathrm{I}}$ given in (24). For applying the principle of typicality, the observer $\mathcal{W}$ is regarded as a quantum system with state space $\mathcal{H}_{\mathcal{W}}$. Then, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{W}}$ is described by a unitary operator $U_{\mathrm{II}}$ such that
$U_{\mathrm{II}}(|\Psi\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle)=\sum_{l=0,1,2}(M^{\mathcal{W}}_{l}|\Psi\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle$ (27)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{F}}$. The unitary operator $U_{\mathrm{II}}$ describes the interaction between the observer $\mathcal{F}$ and the observer $\mathcal{W}$ where both the observers are regarded as a quantum system. The vector $|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{W}}$ is the initial state of the observer $\mathcal{W}$, and $|\Phi^{\mathcal{W}}[l]\rangle\in\mathcal{H}_{\mathcal{W}}$ is a final state of the observer $\mathcal{W}$ for each $l=0,1,2$, with $\langle\Phi^{\mathcal{W}}[l]|\Phi^{\mathcal{W}}[l^{\prime}]\rangle=\delta_{l,l^{\prime}}$. For each $l=0,1$, the state $|\Phi^{\mathcal{W}}[l]\rangle$ indicates that the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $l$.
We denote the set $\{0,1\}\times\{0,1,2\}$ by $\Omega$, and we define a finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
$M_{k,l}:=\delta_{k,l}M^{\mathcal{F}}_{k}.$ (28)
It follows from (23) that
$\sum_{(k,l)\in\Omega}M_{k,l}^{\dagger}M_{k,l}=\sum_{k=0,1}{M^{\mathcal{F}}_{k}}^{\dagger}M^{\mathcal{F}}_{k}=I_{\mathcal{S}},$
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ forms measurement operators. We denote the identity operator acting on $\mathcal{H}_{\mathrm{W}}$ by $I_{\mathcal{W}}$. Then the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
$\begin{split}U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)&:=((I_{\mathcal{S}}\otimes U_{\mathrm{II}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}))(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes\sum_{l=0,1,2}(M^{\mathcal{W}}_{l}|\Phi^{\mathcal{F}}[k]\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[k]\rangle\\ &=\sum_{k=0,1}\,\sum_{l=0,1,2}(M_{k,l}|\Psi\rangle)\otimes|\Phi[k,l]\rangle\\ &=\sum_{(k,l)\in\Omega}(M_{k,l}|\Psi\rangle)\otimes|\Phi[k,l]\rangle\end{split}$ (29)
for each state $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$, where $|\Phi_{\mathrm{init}}\rangle:=|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ and $|\Phi[k,l]\rangle:=|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle$.
9.1 Application of the principle of typicality
The unitary operator $U_{\mathrm{whole}}$ applying to the initial state $|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurement $\mathcal{M}^{\mathcal{F}}$ and the measurement $\mathcal{M}^{\mathcal{W}}$ is infinitely repeated. It follows from (29) that the application of $U_{\mathrm{whole}}$, consisiting of the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ in this order, can be regarded as a single measurement which is described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ and whose all possible outcomes form the set $\Omega$.
Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Thus, according to Definition 6.1, we can see that a world is an infinite sequence over $\Omega$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$, where $P$ is a finite probability space on $\Omega$ such that $P(k,l)$ is the square of the norm of the vector
$(M_{k,l}|+\rangle)\otimes|\Phi[k,l]\rangle$
for every $(k,l)\in\Omega$. Here $\Omega$ is the set of all possible records of the observers $\mathcal{F}$ and $\mathcal{W}$ in the repeated once of the experiments. Let us calculate the explicit form of $P(k,l)$. Then, using (28) and (23), we have that
$P(k,l)=\langle+|M_{k,l}^{\dagger}M_{k,l}|+\rangle\langle\Phi[k,l]|\Phi[k,l]\rangle=\delta_{k,l}\langle+|{M^{\mathcal{F}}_{k}}^{\dagger}M^{\mathcal{F}}_{k}|+\rangle=\delta_{k,l}\left\lvert\langle k|+\rangle\right\rvert^{2}=\frac{\delta_{k,l}}{2}$ (30)
for each $(k,l)\in\Omega$.
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurement described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{W}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{0,1,2\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (30) and Corollary 4.4 that
$\alpha(n)=\beta(n)$ (31)
for every $n\in\mathbb{N}^{+}$. Thus, it follows from Postulate 5 and (29) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{W}$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
$\begin{split}(M_{\alpha(n),\beta(n)}|+\rangle)\otimes|\Phi[\alpha(n),\beta(n)]\rangle&=(M_{\alpha(n),\alpha(n)}|+\rangle)\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\alpha(n)]\rangle\\ &=(M^{\mathcal{F}}_{\alpha(n)}|+\rangle)\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\alpha(n)]\rangle\\ &=\frac{1}{\sqrt{2}}|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\alpha(n)]\rangle,\end{split}$ (32)
up to the normalization factor, where the last equality follows from (23) and (25). Recall that, for each $k=0,1$, the state $|\Phi^{\mathcal{F}}[k]\rangle$ indicates that the observer $\mathcal{F}$ records the value $k$, and the state $|\Phi^{\mathcal{W}}[k]\rangle$ indicates that the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $k$. Thus we have that, in our world $\alpha$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$, the value which the observer $\mathcal{F}$ knows equals the value which the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records, in every repetition of the experiment, independently of $n$, as expected from the point of view of the conventional quantum mechanics.
Let us determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.1, it is easy to check that the states of the system $\mathcal{S}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the states of the system $\mathcal{S}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.1 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the system $\mathcal{S}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\alpha(n)\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (i) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
$|\alpha(n)\rangle$ (33)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$.
On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\Phi^{\mathcal{F}}[\alpha(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
$|\Phi^{\mathcal{F}}[\alpha(n)]\rangle$ (34)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$.
Hence, it follows from (33) and (34) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle$ (35)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. This state (35) is different from the ‘undecided’ state (26) in Case 2. Therefore, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the observer $\mathcal{F}$ surely records the value $\alpha(n)$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in the $n$th repetition of the experiment, in accordance with our common sense.
In addition, it follows from (30) and Theorem 4.1 that $\alpha$ is a Martin-Löf $U$-random infinite binary sequence, where $U$ is a finite probability space on $\{0,1\}$ such that $U(0)=U(1)=1/2$. Namely, $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
$\lim_{n\to\infty}\frac{N_{k}(\alpha\!\!\restriction_{n})}{n}=\frac{1}{2},$
where $N_{k}(\sigma)$ denotes the number of the occurrences of $k$ in $\sigma$ for every $k\in\{0,1\}$ and every $\sigma\in\{0,1\}^{*}$, as expected from the point of view of the conventional quantum mechanics. Thus, in our world $\omega$, for each $k=0,1$ the following (i) and (ii) simultaneously occur in half of the infinite repetitions of the experiment: (i) The observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, and (ii) the observer $\mathcal{F}$ records the value $k$ and the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$, in accordance with our common sense.
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ form PVMs. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle&:=(U_{\mathrm{I}}\otimes I_{\mathcal{W}})(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle,\end{split}$ (36)
where the second and last equalities follow from (24) and (23), respectively. On the other hand, the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$, as we saw above. Thus, it follows from Postulate 7 and (32) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\right)|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ (37)
up to the normalization factor, where
$P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\right):=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\alpha(n)]\rangle\langle\Phi^{\mathcal{F}}[\alpha(n)]\rvert\otimes I_{\mathcal{W}}.$
The vector (37) equals
$\frac{1}{\sqrt{2}}|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$
for every $n\in\mathbb{N}^{+}$, due to (36). Thus, we have the same result as (35), which we obtained based on Postulate 6.
10 Deutsch’s thought experiment
In 1985, modifying the Wigner’s friend paradox described in Section 8, Deutsch [11] proposed a thought experiment, which can, in principle, determine whether the observation by the observer $\mathcal{F}$ can lead to the reduction of state vector of the system $\mathcal{S}$ prepared initially in the state $|+\rangle$. In this paper, we make an analysis of Deutsch’s thought experiment in the framework of quantum mechanics based on the principle of typicality. Actually, in this paper, we study a well-known variant of this thought experiment, which is a simplification of the technique used in Frauchiger and Renner [23]. We call this variant just Deutsch’s thought experiment.
In this section, we review Deutsch’s thought experiment in the terminology of the conventional quantum mechanics. In the setting of Deutsch’s thought experiment, we consider a single qubit system $\mathcal{S}$ and two observers $\mathcal{F}$ and $\mathcal{D}$, as in the original Wigner’s friend paradox described in Section 8. The observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ over the system $\mathcal{S}$ in the exactly same manner as in the Wigner’s friend paradox. However, in Deutsch’s thought experiment [11], the observer $\mathcal{D}$ performs a measurement over the composite system $\mathcal{S}+\mathcal{F}$, and not over the observer $\mathcal{F}$ only, unlike in the case of the original Wigner’s friend paradox. For completeness and clarity, in what follows, we describe Deutsch’s thought experiment from scratch.
Consider a single qubit system $\mathcal{S}$ with state space $\mathcal{H}_{\mathcal{S}}$. An observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ described by the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ over the system $\mathcal{S}$ (according to Postulate 4), where
$M^{\mathcal{F}}_{0}:=\lvert 0\rangle\langle 0\rvert\quad\text{and}\quad M^{\mathcal{F}}_{1}:=\lvert 1\rangle\langle 1\rvert$, (38)
and $|0\rangle$ and $|1\rangle$ form an orthonormal basis of the state space $\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$.
Hereafter, we assume that the observer $\mathcal{F}$ itself is a quantum system with state space $\mathcal{H}_{\mathcal{F}}$. Then we consider an observer $\mathcal{D}$ who is on the outside of the composite system $\mathcal{S}+\mathcal{F}$ consisting of the system $\mathcal{S}$ and the observer $\mathcal{F}$. The observer $\mathcal{D}$ can also be regarded as an environment for the composite system $\mathcal{S}+\mathcal{F}$. Then, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is described by a unitary operator $U_{\mathrm{I}}$ such that
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (39)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. The unitary operator $U_{\mathrm{I}}$ describes the interaction between the system $\mathcal{S}$ and the observer $\mathcal{F}$ as a quantum system. The vector $|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{F}}$ is the initial state of the observer $\mathcal{F}$, and $|\Phi^{\mathcal{F}}[k]\rangle\in\mathcal{H}_{\mathcal{F}}$ is a final state of the observer $\mathcal{F}$ for each $k=0,1$, with $\langle\Phi^{\mathcal{F}}[k]|\Phi^{\mathcal{F}}[k^{\prime}]\rangle=\delta_{k,k^{\prime}}$. For each $k=0,1$, the state $|\Phi^{\mathcal{F}}[k]\rangle$ indicates that the observer $\mathcal{F}$ records the value $k$. The unitary time-evolution (39) is the quantum mechanical description of the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ over the composite system $\mathcal{S}+\mathcal{F}$, from the point of view of the observer $\mathcal{D}$.
We define the state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle:=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$ (40)
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
$\begin{split}M^{\mathcal{D}}_{+}&:=\lvert\Psi^{\mathcal{F}+\mathcal{S}}[+]\rangle\langle\Psi^{\mathcal{F}+\mathcal{S}}[+]\rvert,\\ M^{\mathcal{D}}_{-}&:=\sqrt{I_{\mathcal{S}+\mathcal{F}}-{M^{\mathcal{D}}_{+}}^{\dagger}M^{\mathcal{D}}_{+}}=I_{\mathcal{S}+\mathcal{F}}-M^{\mathcal{D}}_{+},\end{split}$ (41)
and $I_{\mathcal{S}+\mathcal{F}}$ denotes the identity operator acting on $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$.
Now, let us assume that the system $\mathcal{S}$ is initially in the state $|+\rangle$ in the above setting, where
$|+\rangle:=\frac{1}{\sqrt{2}}\left(|0\rangle+|1\rangle\right).$ (42)
Then, we have the following two cases, depending on how we treat the measurement $\mathcal{M}^{\mathcal{F}}$:
Case 1.
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
  1. (i)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|0\rangle$ and $|\Phi^{\mathcal{F}}[0]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle$.
  2. (ii)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|1\rangle$ and $|\Phi^{\mathcal{F}}[1]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle$.
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. Note that, in this Case 1, according to Postulate 4 we can consider the joint probability $P_{\mathrm{c1}}(k,l)$ that the measurement $\mathcal{M}^{\mathcal{F}}$ gives the outcome $k$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $l$ for each $k=0,1$ and $l=+,$-. We then have that $P_{\mathrm{c1}}(k,l)=1/4$ for every $k=0,1$ and $l=+,$-.
Case 2.
According to (39), immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ is
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle.$ (43)
Thus, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the outcome ++ occurs surely in the measurement $\mathcal{M}^{\mathcal{D}}$. Thus, for each $l=+,$-, we can consider the probability $P_{\mathrm{c2}}(l)$ that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $l$, where $P_{\mathrm{c2}}(+)=1$ and $P_{\mathrm{c2}}(-)=0$. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2.
Thus, we have a difference about the measurement results of the measurement $\mathcal{M}^{\mathcal{D}}$, depending on how we treat the measurement $\mathcal{M}^{\mathcal{F}}$, i.e., depending on whether we are based on Postulate 4 (Case 1) or the unitary time-evolution (39) (Case 2). Note that this difference can be testable, in principle, by repeating the experiment sufficiently many times.
11 Analysis of Deutsch’s thought experiment based on the principle of typicality
In this section, we make an analysis of Deutsch’s thought experiment, described in the preceding section, in our framework of quantum mechanics based on Postulates 5 and 7, together with Postulates 1, 2, and 3. We will then see that the measurement $\mathcal{M}^{\mathcal{D}}$ has surely the outcomes ++, even if we fully treat the observer $\mathcal{F}$ as a measuring apparatus.
To proceed this program, first of all we have to implement everything, i.e., the two measurements $\mathcal{M}^{\mathcal{F}}$ and $\mathcal{M}^{\mathcal{D}}$ in the setting of Deutsch’s thought experiment, by unitary time-evolution. The measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is already described by the unitary operator $U_{\mathrm{I}}$ given in (39). For applying the principle of typicality, the observer $\mathcal{D}$ is regarded as a quantum system with state space $\mathcal{H}_{\mathcal{D}}$. Then, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{D}}$ is described by a unitary operator $U_{\mathrm{II}}$ such that
$U_{\mathrm{II}}(|\Psi\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{l=+,-}(M^{\mathcal{D}}_{l}|\Psi\rangle)\otimes|\Phi^{\mathcal{D}}[l]\rangle$ (44)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. The unitary operator $U_{\mathrm{II}}$ describes the interaction between the composite system $\mathcal{S}+\mathcal{F}$ and the observer $\mathcal{D}$ where both are regarded as a quantum system. The vector $|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{D}}$ is the initial state of the observer $\mathcal{D}$, and $|\Phi^{\mathcal{D}}[l]\rangle\in\mathcal{H}_{\mathcal{D}}$ is a final state of the observer $\mathcal{D}$ for each $l=+,$-, with $\langle\Phi^{\mathcal{D}}[l]|\Phi^{\mathcal{D}}[l^{\prime}]\rangle=\delta_{l,l^{\prime}}$. For each $l=+,$-, the state $|\Phi^{\mathcal{D}}[l]\rangle$ indicates that the observer $\mathcal{D}$ records the value $l$.
Let $P_{0}:=\lvert 0\rangle\langle 0\rvert$ and $P_{1}:=\lvert 1\rangle\langle 1\rvert$. Then, from (39), obviously we have that
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}(P_{a}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle$
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. We define $|-\rangle\in\mathcal{H}_{\mathcal{S}}$ by
$|-\rangle:=\frac{1}{\sqrt{2}}\left(|0\rangle-|1\rangle\right).$
Then, it is easy to check that
$\sum_{a=0,1}(P_{a}|+\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle=|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\quad\text{and}\quad\sum_{a=0,1}(P_{a}|-\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\;\perp\;|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle$
hold. Thus, by (39) and (44), we have that
$U_{\mathrm{II}}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|+\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle,$ (45)
$U_{\mathrm{II}}(U_{\mathrm{I}}(|-\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)$ $=\sum_{a=0,1}(P_{a}|-\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle.$ (46)
[…]
$U_{\mathrm{II}}(U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{a=0,1}\sum_{b=+,-}(P_{a}Q_{b}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[a]\rangle\otimes|\Phi^{\mathcal{D}}[b]\rangle$ (47)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. We denote the set $\{0,1\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
$M_{k,l}:=P_{k}Q_{l}.$ (48)
It follows that
$\sum_{(k,l)\in\Omega}M_{k,l}^{\dagger}M_{k,l}=\sum_{b=+,-}\,\sum_{a=0,1}Q_{b}P_{a}Q_{b}=\sum_{b=+,-}Q_{b}=I_{\mathcal{S}},$
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ forms measurement operators. Then, using (47) we see that the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
$U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle):=(U_{\mathrm{II}}\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}))(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)=\sum_{(k,l)\in\Omega}(M_{k,l}|\Psi\rangle)\otimes|\Phi[k,l]\rangle$ (49)
for each state $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$, where $|\Phi_{\mathrm{init}}\rangle:=|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ and $|\Phi[k,l]\rangle:=|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{D}}[l]\rangle$.
11.1 Application of the principle of typicality
The unitary operator $U_{\mathrm{whole}}$ applying to the initial state $|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurement $\mathcal{M}^{\mathcal{F}}$ and the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. It follows from (49) that the application of $U_{\mathrm{whole}}$, consisiting of the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ in this order, can be regarded as a single measurement which is described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ and whose all possible outcomes form the set $\Omega$.
Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Therefore, according to Definition 6.1, we can see that a world is an infinite sequence over $\Omega$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$, where $P$ is a finite probability space on $\Omega$ such that $P(k,l)$ is the square of the norm of the vector
$(M_{k,l}|+\rangle)\otimes|\Phi[k,l]\rangle$
for every $(k,l)\in\Omega$. Here $\Omega$ is the set of all possible records of the observers $\mathcal{F}$ and $\mathcal{D}$ in the repeated once of the experiments. Let us calculate the explicit form of $P(k,l)$. Then, using (48), we have that
$P(k,l)=\langle+|M_{k,l}^{\dagger}M_{k,l}|+\rangle\langle\Phi[k,l]|\Phi[k,l]\rangle=\langle+|Q_{l}P_{k}Q_{l}|+\rangle=\delta_{l,+}\langle+|P_{k}|+\rangle=\frac{\delta_{l,+}}{2}$ (50)
for each $(k,l)\in\Omega$.
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (50) and Corollary 4.4 that $\beta$ consists only of ++, i.e., $\beta=++++++\dotsc\dotsc$. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$\begin{split}(M_{\alpha(n),\beta(n)}|+\rangle)\otimes|\Phi[\alpha(n),\beta(n)]\rangle&=(M_{\alpha(n),+}|+\rangle)\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle\\ &=\frac{1}{\sqrt{2}}|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle,\end{split}$
up to the normalization factor, where the last equality follows from (48). Thus, we have that the observer $\mathcal{D}$ always records the outcome ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment.
In addition, it follows from (50) and Theorem 4.1 that $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
$\lim_{n\to\infty}\frac{N_{k}(\alpha\!\!\restriction_{n})}{n}=\frac{1}{2},$
as expected from the point of view of the conventional quantum mechanics. Thus, in our world $\alpha$, the following (i) and (ii) hold:
  1. (i)
    For each $k=0,1$, the observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$, in half of the infinite repetitions of the experiment.
  2. (ii)
    The observer $\mathcal{D}$ always records the value ++ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment, whichever value the observer $\mathcal{F}$ records in the corresponding repetition.
We can then see that the above situation of (i) and (ii) is different from both Case 1 and Case 2 in Section 10, although the statements of both Case 1 and Case 2 in Section 10 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given, in particular, in the Born rule, Postulate 4 (i).
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. We use $I_{\mathcal{D}}$ to denote the identity operator acting on $\mathcal{H}_{\mathcal{D}}$. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle&:=(U_{\mathrm{I}}\otimes I_{\mathcal{D}})(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (51)
where the second, third, and last equalities follow from (39), (38), and (40), respectively. On the other hand, it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Note also that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. Thus, it follows from Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$P\left(\right)|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ (52)
up to the normalization factor, where $P\left(\right)$ equals the identity operator acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}\otimes\mathcal{H}_{\mathcal{D}}$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$. Hence, it follows from (52) and (51) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle$
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$.
[…]
In the previous section, we have treated the observer $\mathcal{F}$ as a measuring apparatus in the analysis of Deutsch’s thought experiment in our framework of quantum mechanics based on the principle of typicality. What happens if we regard the observer $\mathcal{F}$ as a mere quantum system and not as a measurement apparatus in the analysis of Deutsch’s thought experiment in our framework? We can make an analysis of this case still more easily by the principle of typicality, as follows. We do not need to use either Postulate 6 or Postulate 7 in this case. Note that this case corresponds to Case 2 in Section 10 in the conventional quantum mechanics.
In this case, according to (39), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. Thus, the unitary operator $U_{\mathrm{II}}$ given by (44) applying to the initial state
$(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$
describes the repeated once of the infinite repetition of the measurements (experiments) where the measurement $\mathcal{M}^{\mathcal{D}}$ is infinitely repeated. Let $\Theta:=\{+,-\}$. The application of $U_{\mathrm{II}}$ itself, of course, forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Therefore, according to Definition 6.1, we can see that a world is an infinite sequence over $\Theta$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$, where $R$ is a finite probability space on $\Theta$ such that $R(l)$ is the square of the norm of the vector
$(M^{\mathcal{D}}_{l}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[l]\rangle$
for every $l\in\Theta$. Here $\Theta$ is the set of all possible records of the observer $\mathcal{D}$ in the repeated once of the measurements (experiments). Let us calculate the explicit form of $R(l)$. Recall that
$U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle,$ (53)
from (39), (42), (38), and (40). Thus, using (41) we have that
$R(l)=\langle\Psi^{\mathcal{S}+\mathcal{F}}[+]|{M^{\mathcal{D}}_{l}}^{\dagger}M^{\mathcal{D}}_{l}|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\langle\Phi^{\mathcal{D}}[l]|\Phi^{\mathcal{D}}[l]\rangle=\delta_{l,+}$ (54)
for each $l\in\Theta$.
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observer $\mathcal{D}$ which is being generated by the infinite repetition of the measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{l}\}_{l\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
Then it follows from (54) and Corollary 4.4 that $\omega$ consists only of ++, i.e., $\omega=++++++\dotsc\dotsc$. Thus, it follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the ‘observer’ $\mathcal{F}$ both being only measured, and the real observer $\mathcal{D}$ measuring them, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$(M^{\mathcal{D}}_{\omega(n)}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[\omega(n)]\rangle$ $=(M^{\mathcal{D}}_{+}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{D}}[+]\rangle$
$=|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle,$
where the last equality follows from (41) and (53). It follows that, in our world $\omega$, the observer $\mathcal{D}$ always records the value ++ after the measurement $\mathcal{M}^{\mathcal{D}}$ in every repetition of the experiment. Thus, this result recovers in a refined manner Case 2 in Section 10 which is described in terms of the conventional quantum mechanics. Regarding the results of the measurement $\mathcal{M}^{\mathcal{D}}$ for the observer $\mathcal{D}$, we have the same results as in Section 11, where both treatments are based on the principle of typicality. The important point here is that neither the result of Section 11 nor the result of this section corresponds to Case 1 in Section 10 which is described in terms of the conventional quantum mechanics.
13 Wigner collaborates with Deutsch
Based on the framework developed in Section 7, we can analyze more complicated situations than considered in the previous sections. In this section, as such a example, we consider the combination of Wigner’s friend and Deutsch’s thought experiment. The system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ participate in the experiment. Each of three observers makes their measurements as before but in the order of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$, where the observer $\mathcal{D}$ do nothing to the observer $\mathcal{W}$. We call this combination the Wigner-Deutsch collaboration. In this section, we describe the Wigner-Deutsch collaboration in detail in the terminology of the conventional quantum mechanics from scratch.
Consider a single qubit system $\mathcal{S}$ with state space $\mathcal{H}_{\mathcal{S}}$. An observer $\mathcal{F}$ performs a measurement $\mathcal{M}^{\mathcal{F}}$ described by the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ over the system $\mathcal{S}$ (according to Postulate 4), where
$M^{\mathcal{F}}_{0}:=\lvert 0\rangle\langle 0\rvert\quad\text{and}\quad M^{\mathcal{F}}_{1}:=\lvert 1\rangle\langle 1\rvert$, (55)
and $|0\rangle$ and $|1\rangle$ form an orthonormal basis of the state space $\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$.
Hereafter, we assume that the observer $\mathcal{F}$ itself is a quantum system with state space $\mathcal{H}_{\mathcal{F}}$. Then we consider observers $\mathcal{W}$ and $\mathcal{D}$ who are on the outside of the composite system $\mathcal{S}+\mathcal{F}$ consisting of the system $\mathcal{S}$ and the observer $\mathcal{F}$. The observers $\mathcal{W}$ and $\mathcal{D}$ together can also be regarded as an environment for the composite system $\mathcal{S}+\mathcal{F}$. Then, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is described by a unitary operator $U_{\mathrm{I}}$ such that
$U_{\mathrm{I}}(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (56)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. The unitary operator $U_{\mathrm{I}}$ describes the interaction between the system $\mathcal{S}$ and the observer $\mathcal{F}$ as a quantum system. The vector $|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{F}}$ is the initial state of the observer $\mathcal{F}$, and $|\Phi^{\mathcal{F}}[k]\rangle\in\mathcal{H}_{\mathcal{F}}$ is a final state of the observer $\mathcal{F}$ for each $k=0,1$, with $\langle\Phi^{\mathcal{F}}[k]|\Phi^{\mathcal{F}}[k^{\prime}]\rangle=\delta_{k,k^{\prime}}$. For each $k=0,1$, the state $|\Phi^{\mathcal{F}}[k]\rangle$ indicates that the observer $\mathcal{F}$ records the value $k$. The unitary time-evolution (56) is the quantum mechanical description of the measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ over the composite system $\mathcal{S}+\mathcal{F}$, from the point of view of the observers $\mathcal{W}$ and $\mathcal{D}$.
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
$\begin{split}M^{\mathcal{W}}_{0}&:=\lvert\Phi^{\mathcal{F}}[0]\rangle\langle\Phi^{\mathcal{F}}[0]\rvert,\\ M^{\mathcal{W}}_{1}&:=\lvert\Phi^{\mathcal{F}}[1]\rangle\langle\Phi^{\mathcal{F}}[1]\rvert,\\ M^{\mathcal{W}}_{2}&:=\sqrt{I_{\mathcal{F}}-{M^{\mathcal{W}}_{0}}^{\dagger}M^{\mathcal{W}}_{0}-{M^{\mathcal{W}}_{1}}^{\dagger}M^{\mathcal{W}}_{1}}=I_{\mathcal{F}}-M^{\mathcal{W}}_{0}-M^{\mathcal{W}}_{1},\end{split}$ (57)
and $I_{\mathcal{F}}$ denotes the identity operator acting on $\mathcal{H}_{\mathcal{F}}$.
We define the state $|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle:=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$ (58)
After the measurement $\mathcal{M}^{\mathcal{W}}$ by $\mathcal{W}$ on the observer $\mathcal{F}$, the observer $\mathcal{D}$ performs a measurement $\mathcal{M}^{\mathcal{D}}$ described by the measurement operators $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ over the composite system $\mathcal{S}+\mathcal{F}$ (according to Postulate 4), where
$\begin{split}M^{\mathcal{D}}_{+}&:=\lvert\Psi^{\mathcal{F}+\mathcal{S}}[+]\rangle\langle\Psi^{\mathcal{F}+\mathcal{S}}[+]\rvert,\\ M^{\mathcal{D}}_{-}&:=\sqrt{I_{\mathcal{S}+\mathcal{F}}-{M^{\mathcal{D}}_{+}}^{\dagger}M^{\mathcal{D}}_{+}}=I_{\mathcal{S}+\mathcal{F}}-M^{\mathcal{D}}_{+},\end{split}$ (59)
and $I_{\mathcal{S}+\mathcal{F}}$ denotes the identity operator acting on $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$.
Now, let us assume that the system $\mathcal{S}$ is initially in the state $|+\rangle$ in the above setting, where
$|+\rangle:=\frac{1}{\sqrt{2}}\left(|0\rangle+|1\rangle\right).$ (60)
Then, we have the following two cases, depending on how we treat the measurement $\mathcal{M}^{\mathcal{F}}$:
Case 1.
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
  1. (i)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|0\rangle$ and $|\Phi^{\mathcal{F}}[0]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle$.
  2. (ii)
    The system $\mathcal{S}$ and the observer $\mathcal{F}$ are in the states $|1\rangle$ and $|\Phi^{\mathcal{F}}[1]\rangle$, respectively. Therefore, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle$.
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{W}}$, we see that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome 0 with certainly in the case of the occurrence of the possibility (i), and it gives the outcome $1$ with certainly in the case of the occurrence of the possibility (ii). The measurement $\mathcal{M}^{\mathcal{W}}$ does not change the state of the composite system $\mathcal{S}+\mathcal{F}$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{F}}$. In this Case 1, we can consider the joint probability $P_{\mathrm{c1}}(k,l,m)$ that the measurement $\mathcal{M}^{\mathcal{F}}$ gives the outcome $k$ and then the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $k=0,1$, $l=0,1,2$, and $m=+,$-. We have that $P_{\mathrm{c1}}(k,l,m)=\delta_{k,l}/4$ for every $k=0,1$, $l=0,1,2$, and $m=+,$-. Then the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle$
if the measurement $\mathcal{M}^{\mathcal{D}}$ gives an outcome $m\in\{+,-\}$, where the state $|\Psi^{\mathcal{S}+\mathcal{F}}[-]\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ is defined by
$|\Psi^{\mathcal{S}+\mathcal{F}}[-]\rangle:=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle-|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$
Case 2.
According to (56), immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ is
$|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle.$
Then, applying Postulate 4 to the subsequent measurement $\mathcal{M}^{\mathcal{W}}$, we see that the measurement $\mathcal{M}^{\mathcal{W}}$ results in either the following two possibilities (i) or (ii):
  1. (i)
    The measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome 0, and the post-measurement state of the composite system $\mathcal{S}+\mathcal{F}$ is $|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle$.
  2. (ii)
    The measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $1$, and the post-measurement state of the composite system $\mathcal{S}+\mathcal{F}$ is $|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle$.
Here, each possibility occurs with the same probability $1/2$. Then, applying Postulate 4 to the further subsequent measurement $\mathcal{M}^{\mathcal{D}}$, we see that the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcomes ++ with probability $1/2$, whichever of the two possibilities (i) or (ii) above occurs in the measurement $\mathcal{M}^{\mathcal{W}}$. Thus, in this Case 2, we can consider the joint probability $P_{\mathrm{c2}}(l,m)$ that the measurement $\mathcal{M}^{\mathcal{W}}$ gives the outcome $l$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ gives the outcome $m$ for each $l=0,1,2$ and $m=+,$-. We have that $P_{\mathrm{c2}}(l,m)=1/4$ for every $l=0,1$ and $m=+,$-, and $P_{\mathrm{c2}}(2,m)=0$ for every $m=+,$-. On the other hand, the notion of the probability that the measurement $\mathcal{M}^{\mathcal{F}}$ gives a certain outcome is meaningless in this Case 2. In the same manner as in Case 1, the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle$
if the measurement $\mathcal{M}^{\mathcal{D}}$ gives an outcome $m\in\{+,-\}$.
Thus, if we ignore the measurement results of the measurement $\mathcal{M}^{\mathcal{F}}$ in Case 1, there is no difference between Case 1 and Case 2 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$. Actually, we have that
$\sum_{k=0,1}P_{\mathrm{c1}}(k,l,m)=\frac{1}{4}=P_{\mathrm{c2}}(l,m)$
for every $l=0,1$ and $m=+,$-, and that
$\sum_{k=0,1}P_{\mathrm{c1}}(k,2,m)=0=P_{\mathrm{c2}}(2,m)$
for every $m=+,$-.
14 Analysis of the Wigner-Deutsch collaboration based on the principle of typicality
In this section, we make an analysis of the Wigner-Deutsch collaboration described in the preceding section, in our framework of quantum mechanics based on Postulates 5, 6, and 7, together with Postulates 1, 2, and 3.
To proceed this program, first of all we have to implement everything, i.e., the three measurements $\mathcal{M}^{\mathcal{F}}$, $\mathcal{M}^{\mathcal{W}}$, and $\mathcal{M}^{\mathcal{D}}$ in the setting of the Wigner-Deutsch collaboration, by unitary time-evolution. The measurement process of the measurement $\mathcal{M}^{\mathcal{F}}$ is already described by the unitary operator $U_{\mathrm{I}}$ given in (56). For applying the principle of typicality, the observers $\mathcal{W}$ and $\mathcal{D}$ are regarded as quantum systems with state spaces $\mathcal{H}_{\mathcal{W}}$ and $\mathcal{H}_{\mathcal{D}}$, respectively. Then, on the one hand, according to (5), the measurement process of the measurement $\mathcal{M}^{\mathcal{W}}$ is described by a unitary operator $U_{\mathrm{II}}$ such that
$U_{\mathrm{II}}(|\Psi\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle)=\sum_{l=0,1,2}(M^{\mathcal{W}}_{l}|\Psi\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle$ (61)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{F}}$. The unitary operator $U_{\mathrm{II}}$ describes the interaction between the observer $\mathcal{F}$ and the observer $\mathcal{W}$ where both the observers are regarded as a quantum system. The vector $|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{W}}$ is the initial state of the observer $\mathcal{W}$, and $|\Phi^{\mathcal{W}}[l]\rangle\in\mathcal{H}_{\mathcal{W}}$ is a final state of the observer $\mathcal{W}$ for each $l=0,1,2$, with $\langle\Phi^{\mathcal{W}}[l]|\Phi^{\mathcal{W}}[l^{\prime}]\rangle=\delta_{l,l^{\prime}}$. For each $l=0,1$, the state $|\Phi^{\mathcal{W}}[l]\rangle$ indicates that the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $l$. On the other hand, according to (5) again, the measurement process of the measurement $\mathcal{M}^{\mathcal{D}}$ is described by a unitary operator $U_{\mathrm{III}}$ such that
$U_{\mathrm{III}}(|\Psi\rangle\otimes|\Phi\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)=\sum_{m=+,-}(M^{\mathcal{D}}_{m}|\Psi\rangle)\otimes|\Phi\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle$ (62)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ and $|\Phi\rangle\in\mathcal{H}_{\mathcal{W}}$. The unitary operator $U_{\mathrm{III}}$ describes the interaction between the composite system $\mathcal{S}+\mathcal{F}$ and the observer $\mathcal{D}$ where both are regarded as a quantum system. The vector $|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\in\mathcal{H}_{\mathcal{D}}$ is the initial state of the observer $\mathcal{D}$, and $|\Phi^{\mathcal{D}}[m]\rangle\in\mathcal{H}_{\mathcal{D}}$ is a final state of the observer $\mathcal{D}$ for each $m=+,$-, with $\langle\Phi^{\mathcal{D}}[m]|\Phi^{\mathcal{D}}[m^{\prime}]\rangle=\delta_{m,m^{\prime}}$. For each $m=+,$-, the state $|\Phi^{\mathcal{D}}[m]\rangle$ indicates that the observer $\mathcal{D}$ records the value $m$.
Now, for each $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ to the state $|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$, consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{W}$, result in the state:
$\begin{split}&((I_{\mathcal{S}}\otimes U_{\mathrm{II}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}))(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle)\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes\sum_{l=0,1,2}(M^{\mathcal{W}}_{l}|\Phi^{\mathcal{F}}[k]\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[k]\rangle,\end{split}$ (63)
where $I_{\mathcal{S}}$ and $I_{\mathcal{W}}$ denote the identity operators acting on $\mathcal{H}_{\mathcal{S}}$ and $\mathcal{H}_{\mathcal{W}}$, respectively. Thus, it follows from (62) that
$\begin{split}&(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|0\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\frac{1}{2}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle\otimes|\Phi^{\mathcal{W}}[0]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle\otimes|\Phi^{\mathcal{W}}[0]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle\\ &\quad+|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle\otimes|\Phi^{\mathcal{W}}[0]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle-|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle\otimes|\Phi^{\mathcal{W}}[0]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle),\end{split}$ (64)
and
$\begin{split}&(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|1\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\frac{1}{2}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle\otimes|\Phi^{\mathcal{W}}[1]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle\otimes|\Phi^{\mathcal{W}}[1]\rangle\otimes|\Phi^{\mathcal{D}}[+]\rangle\\ &\quad-|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle\otimes|\Phi^{\mathcal{W}}[1]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle\otimes|\Phi^{\mathcal{W}}[1]\rangle\otimes|\Phi^{\mathcal{D}}[-]\rangle),\end{split}$ (65)
where $I_{\mathcal{D}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{D}}$. We define $|-\rangle\in\mathcal{H}_{\mathcal{S}}$ by
$|-\rangle:=\frac{1}{\sqrt{2}}\left(|0\rangle-|1\rangle\right).$
[…]
$\begin{split}&(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|\Psi\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{k=0,1}\sum_{l=0,1}\sum_{m=+,-}(P_{k}Q_{m}P_{l}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle\end{split}$ (66)
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. We denote the set $\{0,1\}\times\{0,1,2\}\times\{+,-\}$ by $\Omega$, and we define a finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}$ by
$M_{k,l,m}:=P_{k}Q_{m}P_{l}$ (67)
if $l\neq 2$ and $M_{k,l,m}:=0$ otherwise. It follows that
$\sum_{(k,l,m)\in\Omega}M_{k,l,m}^{\dagger}M_{k,l,m}=\sum_{l=0,1}\,\sum_{m=+,-}\,\sum_{k=0,1}P_{l}Q_{m}P_{k}Q_{m}P_{l}=\sum_{l=0,1}\,\sum_{m=+,-}P_{l}Q_{m}P_{l}=\sum_{l=0,1}P_{l}=I_{\mathcal{S}}.$
Thus, the finite collection $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ forms measurement operators. Then, using (66) we see that, as a whole, the sequential applications of $U_{\mathrm{I}}$, $U_{\mathrm{II}}$, and $U_{\mathrm{III}}$ result in:
$\begin{split}U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)&:=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|\Psi\rangle\otimes|\Phi_{\mathrm{init}}\rangle)\\ &=\sum_{(k,l,m)\in\Omega}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi[k,l,m]\rangle\end{split}$ (68)
for each state $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ of the system $\mathcal{S}$, where $|\Phi_{\mathrm{init}}\rangle:=|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ and $|\Phi[k,l,m]\rangle:=|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle$.
14.1 Application of the principle of typicality
The unitary operator $U_{\mathrm{whole}}$ applying to the initial state $|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle$ describes the repeated once of the infinite repetition of the measurements where the succession of the measurements $\mathcal{M}^{\mathcal{F}}$, $\mathcal{M}^{\mathcal{W}}$, and $\mathcal{M}^{\mathcal{D}}$ in this order is infinitely repeated. It follows from (68) that the application of $U_{\mathrm{whole}}$, consisting of the sequential applications of $U_{\mathrm{I}}$, $U_{\mathrm{II}}$, and $U_{\mathrm{III}}$ in this order, can be regarded as a single measurement which is described by the measurement operators $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ and whose all possible outcomes form the set $\Omega$.
Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Therefore, according to Definition 6.1, we can see that a world is an infinite sequence over $\Omega$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$, where $P$ is a finite probability space on $\Omega$ such that $P(k,l,m)$ is the square of the norm of the vector
$(M_{k,l,m}|+\rangle)\otimes|\Phi[k,l,m]\rangle$
for every $(k,l,m)\in\Omega$. Here $\Omega$ is the set of all possible records of the observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ in the repeated once of the experiments. Let us calculate the explicit form of $P(k,l,m)$. We use $\Omega_{c}$ to denote the set $\{0,1\}\times\{0,1\}\times\{+,-\}$, which is a proper subset of $\Omega$. First, using (67) we can see that
$(M_{k,l,m}|+\rangle)\otimes|\Phi[k,l,m]\rangle=0$ (69)
for every $(k,l,m)\in\Omega\setminus\Omega_{c}$, and
$(M_{k,l,m}|+\rangle)\otimes|\Phi[k,l,m]\rangle=\frac{f(k,l,m)}{2\sqrt{2}}|k\rangle\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle$ (70)
for every $(k,l,m)\in\Omega_{c}$, where $f$ denotes a function from $\Omega_{c}$ to $\{1,-1\}$ such that f(k,l,m)=1f(k,l,m)=-1 iff $k\neq l$ and $m=$-. Thus, using (69) and (70), we have that
$P(k,l,m)=0$ (71)
for every $(k,l,m)\in\Omega\setminus\Omega_{c}$, and
$P(k,l,m)=\frac{1}{8}$ (72)
for every $(k,l,m)\in\Omega_{c}$.
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l,m}\}_{(k,l,m)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$, $\beta$, and $\gamma$ to denote the infinite sequences over $\{0,1\}$, $\{0,1,2\}$, and $\{+,-\}$, respectively, such that $(\alpha(n),\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (71) and Corollary 4.4 that
$\omega(n)\in\Omega_{c}$
for every $n\in\mathbb{N}^{+}$, and therefore
$\beta(n)\neq 2$
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$\begin{split}&(M_{\alpha(n),\beta(n),\gamma(n)}|+\rangle)\otimes|\Phi[\alpha(n),\beta(n),\gamma(n)]\rangle\\ &=\frac{f(\alpha(n),\beta(n),\gamma(n))}{2\sqrt{2}}|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}[\gamma(n)]\rangle\end{split}$ (73)
up to the normalization factor, where the equality follows from (70). Since $\omega$ is Martin-Löf $P$-random, it follows from (72) and Theorem 4.2 that for every $(k,l,m)\in\Omega_{c}$ it holds that
$\lim_{n\to\infty}\frac{N_{k,l,m}(\omega\!\!\restriction_{n})}{n}=\frac{1}{8},$ (74)
where $N_{k,l,m}(\sigma)$ denotes the number of the occurrences of $(k,l,m)$ in $\sigma$ for every $(k,l,m)\in\Omega_{c}$ and every $\sigma\in\Omega_{c}^{*}$. Thus, in our world $\omega$, the following holds for each $(k,l,m)\in\Omega_{c}$: In one-eighth of the infinite repetitions of the experiment, the state of $\mathcal{S}$ is $|k\rangle$ and the observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$ record the values $k$, $l$, and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that this situation is different from both Case 1 and Case 2 in Section 13, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given, in particular, in the Born rule, Postulate 4 (i).
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ is confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (73), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is
$|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ (75)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$.
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$, $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$, $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. On the one hand, applying $(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}})$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle&:=((I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[k]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}\sum_{k=0,1}|k\rangle\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[k]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (76)
where the second and last equalities follow from (63) and (55), respectively. On the other hand, the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. But it is easy to see that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (73) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ (77)
up to the normalization factor, where
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
The vector (77) equals
$\frac{1}{\sqrt{2}}|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ (78)
for every $n\in\mathbb{N}^{+}$, due to (76). Hence, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
$|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ (79)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$.
Now, we can apply Postulate 7 recursively, based on (79). Namely, we can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Recall that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ form PVMs. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments where the observer $\mathcal{D}$ is ignored results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle&:=(U_{\mathrm{I}}\otimes I_{\mathcal{W}})(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle)=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle,\end{split}$ (80)
where the second and last equalities follow from (56) and (55), respectively. On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from Postulate 7 and (79) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the two observers $\mathcal{F}$ and $\mathcal{W}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$P\left(|\Phi^{\mathcal{F}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ (81)
up to the normalization factor, where
$P\left(|\Phi^{\mathcal{F}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\beta(n)]\rangle\langle\Phi^{\mathcal{F}}[\beta(n)]\rvert\otimes I_{\mathcal{W}}.$
The vector (81) equals
$\frac{1}{\sqrt{2}}|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$
for every $n\in\mathbb{N}^{+}$, due to (80). Thus, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle$ (82)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Hence, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the observer $\mathcal{F}$ surely records the value $\beta(n)$ regarding the system $\mathcal{S}$ whose state is $|\beta(n)\rangle$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in the $n$th repetition of the experiment, in accordance with our common sense. Note that we can also derive the result (82) from (79) using Postulate 6 in the same manner as we derived the result (35) in Section 9.1, instead of using Postulate 7.
It follows from (74) that
$\begin{split}\lim_{n\to\infty}\frac{N_{l}(\beta\!\!\restriction_{n})}{n}&=\lim_{n\to\infty}\frac{1}{n}\sum_{k=0,1}\sum_{m=+,-}N_{k,l,m}(\omega\!\!\restriction_{n})=\sum_{k=0,1}\sum_{m=+,-}\lim_{n\to\infty}\frac{N_{k,l,m}(\omega\!\!\restriction_{n})}{n}\\ &=\sum_{k=0,1}\sum_{m=+,-}\frac{1}{8}=\frac{1}{2}\end{split}$
for each $l\in\{0,1\}$, where $N_{l}(\sigma)$ denotes the number of the occurrences of $l$ in $\sigma$ for every $l\in\{0,1\}$ and every $\sigma\in\{0,1\}^{*}$. Thus, the states (82) and (79) imply that, in our world $\omega$, the following (i) and (ii) hold for each $l\in\{0,1\}$ and in half of the infinite repetitions of the experiment:
  1. (i)
    The state of $\mathcal{S}$ is $|l\rangle$ and the observer $\mathcal{F}$ records this value $l$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$.
  2. (ii)
    The state of $\mathcal{S}$ is $|l\rangle$ and both the observers $\mathcal{F}$ and $\mathcal{W}$ record this value $l$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$.
We can then see that the this situation regarding $\mathcal{M}^{\mathcal{F}}$ and $\mathcal{M}^{\mathcal{W}}$ coincides with Case 1 in Section 13, although the statement of Case 1 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
Remark 14.1 (The Validity I of Postulate 7).
Let us consider the validity of Postulate 7. In Postulate 7, when acted on the vector $|\Psi^{n-1}_{\mathrm{Total}}\rangle$, the projector
$P(|\Phi^{i_{1}}[m_{i_{1}}]\rangle,\dots,|\Phi^{i_{k}}[m_{i_{k}}]\rangle)$ (83)
determines the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ in a manner that the apparatuses $\mathcal{A}_{i_{1}},\dots,\mathcal{A}_{i_{k}}$ are in the states $|\Phi^{i_{1}}\rangle,\dots,|\Phi^{i_{k}}\rangle$, respectively, immediately after the measurement $\mathcal{M}^{n-1}$. Recall that these apparatuses $\mathcal{A}_{i_{1}},\dots,\mathcal{A}_{i_{k}}$ forms precisely the set of apparatuses among $\mathcal{A}_{1},\dots,\mathcal{A}_{n-1}$ whose final states are unchanged after the measurement by the apparatus $\mathcal{A}_{n}$. If we could replace the projector $P(|\Phi^{i_{1}}[m_{i_{1}}]\rangle,\dots,|\Phi^{i_{k}}[m_{i_{k}}]\rangle)$ by
$P(|\Phi^{1}[m_{1}]\rangle,\dots,|\Phi^{n-1}[m_{n-1}]\rangle)$ (84)
in Postulate 7, we could determine all states of the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n-1}$ immediately after the measurement $\mathcal{M}^{n-1}$. However, this is impossible, as we will see in what follows.
In the Wigner-Deutsch collaboration, the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by (77) in our world $\omega$. This is because the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ while the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Consequently, the projector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (83) in this case. In contrast, the projector $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (84) in this case. Let us calculate $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$. Then using (76) we have that
$P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle=\frac{\delta_{\alpha(n),\beta(n)}}{\sqrt{2}}|\alpha(n)\rangle\otimes|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ (85)
for every $n\in\mathbb{N}^{+}$. However, it follows from (74) that $\alpha(n)\neq\beta(n)$ for half of $n\in\mathbb{N}^{+}$, since
$\lim_{n\to\infty}\frac{1}{n}\sum_{k=0,1}\sum_{m=+,-}N_{k,1-k,m}(\omega\!\!\restriction_{n})=\sum_{k=0,1}\sum_{m=+,-}\lim_{n\to\infty}\frac{N_{k,1-k,m}(\omega\!\!\restriction_{n})}{n}=\sum_{k=0,1}\sum_{m=+,-}\frac{1}{8}=\frac{1}{2}.$
Therefore, the vector (85) becomes 0 for such infinitely many $n$. Hence, the vector (85) does not become the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, since the state vector must not be a zero-vector. In contrast, the vector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$, which the original Postulate 7 requires to be the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, is non-zero for all $n\in\mathbb{N}^{+}$, since it equals (78) for all $n\in\mathbb{N}^{+}$. This observation supports the validity of Postulate 7. ∎
Remark 14.2 (The Validity II of Postulate 7).
Let us consider the validity of Postulate 7 further. Postulate 7 can be applied on the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. We will demonstrate the necessity of this premise, based on the Wigner-Deutsch collaboration. We call a variant of Postulate 7 where this premise is eliminated the naive Postulate 7. We will show that the naive Postulate 7 is too strong and leads to a contradiction.
On the one hand, using the naive Postulate 6 we can derive the result (82) from (73) in the same manner as we derived the result (82) from (73) above using Postulate 7. That is, we have that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle$ (86)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$.
On the other hand, in the Wigner-Deutsch collaboration we can regard the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ as a single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$, where we regard the combination of the observers $\mathcal{W}$ and $\mathcal{D}$ as a single observer $\mathcal{W}+\mathcal{D}$. Actually, the measurement process of the measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$ is described by a unitary operator $U_{\mathcal{W}+\mathcal{D}}$ defined by
$U_{\mathcal{W}+\mathcal{D}}:=U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}).$
We denote the set $\{0,1,2\}\times\{+,-\}$ by $\Theta$, and we define a finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ by
$M^{\mathcal{W}+\mathcal{D}}_{l,m}:=M^{\mathcal{D}}_{m}\circ(I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l}).$
Using (61) and (62) we have that
$U_{\mathcal{W}+\mathcal{D}}(|\Psi\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ =\sum_{(l,m)\in\Theta}(M^{\mathcal{W}+\mathcal{D}}_{l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle$
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$, in the form corresponding to (9). It is then easy to see that the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ forms measurement operators.
Now, we can use the naive Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, under the situation that the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments where the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$ results in a state $|\Psi^{\mathcal{F}(\mathcal{W}+\mathcal{D})}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{F}(\mathcal{W}+\mathcal{D})}_{\mathrm{Total}}\rangle&:=(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}})(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{k=0,1}(M^{\mathcal{F}}_{k}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (87)
where the second and last equalities follow from (56) and (55), respectively. On the other hand, it is easy to check that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{W}+\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from the naive Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$P\left(\right)|\Psi^{\mathcal{F}(\mathcal{W}+\mathcal{D})}_{\mathrm{Total}}\rangle$ (88)
up to the normalization factor, where $P\left(\right)$ equals the identity operator acting on the whole state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}\otimes\mathcal{H}_{\mathcal{W}}\otimes\mathcal{H}_{\mathcal{D}}$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Hence, it follows from (88) and (87) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
$\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)$
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. However, this contradicts the result (86), which implies that the naive Postulate 7 is too strong to be consistent.
In contrast, we can avoid this contradiction for Postulate 7, since the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ do not form a PVM and therefore the above argument does not work for Postulate 7. Hence, in Postulate 7, the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$ is necessary. ∎
[…]
In the previous section, we have treated the observer $\mathcal{F}$ as a measuring apparatus in the analysis of the Wigner-Deutsch collaboration in our framework of quantum mechanics based on the principle of typicality. What happens if we regard the observer $\mathcal{F}$ as a mere quantum system and not as a measurement apparatus in the analysis of the Wigner-Deutsch collaboration in our framework? We can make an analysis of this case more easily by the principle of typicality, as follows. We will see that this case corresponds to Case 2 in Section 13, described in the terminology of the conventional quantum mechanics.
In this case, according to (56), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. The succession of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ performed over the composite system $\mathcal{S}+\mathcal{F}$ in this state forms the repeated once of the infinite repetition of measurement. Thus, for applying the principle of typicality, we have to implement the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ by unitary time-evolution. This is already done in $U_{\mathrm{II}}$ of (61) and $U_{\mathrm{III}}$ of (62).
Now, it follows from (61) and (62) that
$\begin{split}&(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}))(|\Psi\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=U_{\mathrm{III}}\sum_{l=0,1,2}((I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})|\Psi\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=\sum_{l=0,1,2}\,\sum_{m=+,-}((M^{\mathcal{D}}_{m}\circ(I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l}))|\Psi\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle\end{split}$ (89)
for each $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. We denote the set $\{0,1,2\}\times\{+,-\}$ by $\Theta$, and we define a finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
$M^{\mathcal{W}+\mathcal{D}}_{l,m}:=M^{\mathcal{D}}_{m}\circ(I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l}).$ (90)
It follows from (59) and (57) that
$\sum_{(l,m)\in\Theta}{M^{\mathcal{W}+\mathcal{D}}_{l,m}}^{\dagger}M^{\mathcal{W}+\mathcal{D}}_{l,m}$ $=\sum_{l=0,1,2}(I_{\mathcal{S}}\otimes{M^{\mathcal{W}}_{l}}^{\dagger})\left[\sum_{m=+,-}{M^{\mathcal{D}}_{m}}^{\dagger}M^{\mathcal{D}}_{m}\right](I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})$
$=I_{\mathcal{S}}\otimes\sum_{l=0,1,2}{M^{\mathcal{W}}_{l}}^{\dagger}M^{\mathcal{W}}_{l}=I_{\mathcal{S}+\mathcal{F}}.$
Thus, the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ forms measurement operators. Then, using (89) we see that, as a whole, the sequential applications of $U_{\mathrm{II}}$ and $U_{\mathrm{III}}$ to the state $|\Psi\rangle\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ result in the state:
$\begin{split}U_{\mathrm{whole}}(|\Psi\rangle\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)&:=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}))(|\Psi\rangle\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{(l,m)\in\Theta}(M^{\mathcal{W}+\mathcal{D}}_{l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle\end{split}$ (91)
for each state $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$, where $|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle:=|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ and $|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle:=|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle$.
In this case, according to (56), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. Thus, the unitary operator $U_{\mathrm{whole}}$ given by (91) applying to the initial state
$(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$
describes the repeated once of the infinite repetition of the measurements (experiments) where the succession of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ in this order is infinitely repeated. It follows from (91) that the application of $U_{\mathrm{whole}}$ forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
Hence, we can apply Definition 6.1 to this scenario of the setting of measurements. Therefore, according to Definition 6.1, we can see that a world is an infinite sequence over $\Theta$ and the probability measure induced by the probability measure representation for the prefixes of worlds is a Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$, where $R$ is a finite probability space on $\Theta$ such that $R(l,m)$ is the square of the norm of the vector
$(M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle$
for every $(l,m)\in\Theta$. Here $\Theta$ is the set of all possible records of the observers $\mathcal{W}$ and $\mathcal{D}$ in the repeated once of the measurements (experiments). Let us calculate the explicit form of $R(l,m)$. First, it follows from (91) and (68) that
$\begin{split}&\sum_{(l,m)\in\Theta}(M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}))((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle)\\ &=\sum_{(k,l,m)\in\Omega}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi[k,l,m]\rangle\\ &=\sum_{(l,m)\in\Theta}\left\{\sum_{k=0,1}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\right\}\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle.\end{split}$
Thus, we have that
$M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))=\sum_{k=0,1}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (92)
for every $(l,m)\in\Theta$. We use $\Theta_{c}$ to denote the set $\{0,1\}\times\{+,-\}$, which is a proper subset of $\Theta$. On the one hand, using (92) and (67), we have that
$R(l,m)=0$ (93)
for every $(l,m)\in\Theta\setminus\Theta_{c}$. On the other hand, for each $(l,m)\in\Theta_{c}$, using (92) and (70) we have that
$\begin{split}&(M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle\\ &=\sum_{k=0,1}(M_{k,l,m}|+\rangle)\otimes|\Phi[k,l,m]\rangle\\ &=\sum_{k=0,1}\frac{f(k,l,m)}{2\sqrt{2}}|k\rangle\otimes|\Phi^{\mathcal{F}}[k]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle\\ &=\frac{g(l,m)}{2}|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}[m]\rangle,\end{split}$ (94)
where $g$ denotes a function from $\Theta_{c}$ to $\{1,-1\}$ such that g(l,m)=1g(l,m)=-1 iff $l=1$ and $m=$-. Hence, using (94) we have that
$R(l,m)=\frac{\left\lvert g(l,m)\right\rvert^{2}}{4}\langle\Psi^{\mathcal{S}+\mathcal{F}}[m]|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle=\frac{1}{4}$ (95)
for each $(l,m)\in\Theta_{c}$.
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observers $\mathcal{W}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
We use $\beta$ and $\gamma$ to denote the infinite sequences over $\{0,1,2\}$ and $\{+,-\}$, respectively, such that $(\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (93) and Corollary 4.4 that
$\omega(n)\in\Theta_{c}$
for every $n\in\mathbb{N}^{+}$, and therefore
$\beta(n)\neq 2$ (96)
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
$\begin{split}&(M^{\mathcal{W}+\mathcal{D}}_{\beta(n),\gamma(n)}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[\beta(n),\gamma(n)]\rangle\\ &=\frac{g(\beta(n),\gamma(n))}{2}|\Psi^{\mathcal{S}+\mathcal{F}}[\gamma(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}[\gamma(n)]\rangle\end{split}$ (97)
up to the normalization factor, where the equality follows from (94). Since $\omega$ is Martin-Löf $R$-random, it follows from (95) and Theorem 4.2 that for every $(l,m)\in\Theta_{c}$ it holds that
$\lim_{n\to\infty}\frac{N_{l,m}(\omega\!\!\restriction_{n})}{n}=\frac{1}{4},$
where $N_{l,m}(\sigma)$ denotes the number of the occurrences of $(l,m)$ in $\sigma$ for every $(l,m)\in\Theta_{c}$ and every $\sigma\in\Theta_{c}^{*}$. Thus, in our world $\omega$, the following holds for each $(l,m)\in\Theta_{c}$: In one-fourth of the infinite repetitions of the experiment, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle$ and the observers $\mathcal{W}$ and $\mathcal{D}$ record the values $l$ and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that the this situation coincides with that of both Case 1 and Case 2 in Section 13 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ is confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (97), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is
$|\Phi^{\mathcal{W}}[\beta(n)]\rangle$
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$.
We can use Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ and $\{M^{\mathcal{D}}_{+},M^{\mathcal{D}}_{-}\}$ form PVMs. On the one hand, applying $I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}$ to the initial state $(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ given by
$\begin{split}|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle&:=(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=\sum_{l=0,1,2}((I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle)\otimes|\Phi^{\mathcal{W}}[l]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$ (98)
where the equality follows from (56), (55), (58), and (61). On the other hand, the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Thus, it follows from Postulate 7 and (97) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$ (99)
up to the normalization factor, where
$P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right):=I_{\mathcal{S}}\otimes I_{\mathcal{F}}\otimes\lvert\Phi^{\mathcal{W}}[\beta(n)]\rangle\langle\Phi^{\mathcal{W}}[\beta(n)]\rvert\otimes I_{\mathcal{D}}.$
The vector (99) equals
$\begin{split}&((I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{\beta(n)})|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle)\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle\\ &=\frac{1}{\sqrt{2}}|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle,\end{split}$
for every $n\in\mathbb{N}^{+}$, due to (98), (57), and (58). Hence, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by
$|\beta(n)\rangle\otimes|\Phi^{\mathcal{F}}[\beta(n)]\rangle\otimes|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ (100)
in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$.
In addition, it follows from Theorem 4.1, (95), and (96) that $\beta$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $l\in\{0,1\}$ it holds that
$\lim_{n\to\infty}\frac{N_{l}(\beta\!\!\restriction_{n})}{n}=\frac{1}{2},$
where $N_{l}(\sigma)$ denotes the number of the occurrences of $l$ in $\sigma$ for every $l\in\{0,1\}$ and every $\sigma\in\{0,1\}^{*}$. Thus, the state (100) implies that, in our world $\omega$, the following holds for each $l\in\{0,1\}$ and in half of the infinite repetitions of the experiment: The state of $\mathcal{S}$ is $|l\rangle$ and both the observers $\mathcal{F}$ and $\mathcal{W}$ record this value $l$, immediately after the measurement $\mathcal{M}^{\mathcal{W}}$.
In our framework of quantum mechanics based on the principle of typicality, we have investigated the Wigner-Deutsch collaboration in the case where the observer $\mathcal{F}$ is treated as a mere quantum system and not as a measurement apparatus. Based on the above analysis, we can see that this case recovers in a refined manner Case 2 in Section 13, which is described in the terminology of the conventional quantum mechanics in a vague manner. Regarding the results of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$, this case coincides with Case 1 in Section 13 as well, which is also described in terms of the conventional quantum mechanics in a vague manner.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 22K03409.